### v1

# Privacy Technologies for Financial Intelligence

Abstract
Financial crimes like terrorism financing and money laundering can have real impacts on society, including the abuse and mismanagement of public funds, increase in societal problems such as drug trafficking and illicit gambling with attendant economic costs, and loss of innocent lives in the case of terrorism activities. Complex financial crimes can be hard to detect primarily because data related to different pieces of the overall puzzle is usually distributed across a network of financial institutions, regulators, and law-enforcement agencies and they cannot be easily shared due to privacy constraints. Recent advances in Privacy-Preserving Data Matching and Machine Learning provide an opportunity for regulators and the financial industry to come together to solve the risk-discovery problem with technology. This paper provides a survey of the financial intelligence landscape and where opportunities lie for privacy technologies to improve the state-of-the-art in financial-crime detection.
1 Introduction
Our economy depends on a well functioning banking and financial system that has the trust of the general population. However in recent years, the reputation of some major banks has been severely damaged as criminal elements have exploited systemic weaknesses to launder huge sums of money, including transfers to terrorist organisations [Lyn22, LF23]. Besides reputational damage to financial institutions, complex financial crimes like terrorism financing and money laundering can have real direct impacts on society too, including
  • abuse and mismanagement of public funds;
  • increase in societal problems like drug trafficking, illicit trade and gambling, people smuggling, with attendant economic costs; and
  • loss of innocent lives in the case of terrorism activities.
Complex financial crimes can be hard to detect primarily because data related to different pieces of the overall puzzle are usually distributed across a network of financial institutions, regulators, and law-enforcement agencies [LCSH19]. The problem is also rapidly increasing in complexity because of new emerging value-transfer platforms like crypto-currencies and real-time settlement systems.
In early 2016, AUSTRAC, Australia’s Financial Intelligence Agency, set up the Fintel Alliance [^1] https://www.austrac.gov.au/partners/fintel-alliance, an organisation that includes AUSTRAC and the major financial institutions and law-enforcement agencies, to improve public-private collaboration in combating financial crimes. Since then, financial intelligence agencies around the world have started building public-private data-sharing platforms to share and disrupt criminal entities engaging in financial crimes, including the Joint Money Laundering Intelligence Task Force (JMLIT) [^2] https://www.nationalcrimeagency.gov.uk/what-we-do/national-economic-crime-centre in UK, and the Financial Crime Enforcement Network Exchange [^3] https://www.fincen.gov/resources/financial-crime-enforcement-network-exchange in the United States. These public-private platforms are highly effective against known threats.
However, their discovery functions to find as-yet unknown risks in the financial system are still not as strong as they can be, primarily because of privacy constraints around large-scale data matching and analytics between organisations. Recent advances in Privacy-Preserving Data Matching and Machine Learning provide an opportunity for regulators and the financial industry to come together to solve the risk-discovery problem with technology. This paper provides a survey of the financial intelligence landscape and where opportunities lie for privacy technologies to improve the state-of-the-art in financial-crime detection. Our paper complements the material covered in [Max20], which explored global case studies of public and private sectors employing cryptographic-based privacy technologies to foster collaborative financial intelligence. [^4] https://www.future-fis.com/the-pet-project.html
2 Financial Intelligence
Financial intelligence is a highly specialised field that is not well understood by many. We provide an introduction to the financial intelligence ecosystem in this section, including a description of the current state, some existing limitations and opportunities.
2.1 The Financial Intelligence Ecosystem
Financial Intelligence Units (FIUs) around the world are governed by anti-money laundering and counter-terrorism financing (AML/CTF) regimes that are quite similar in nature [Mad11, Gil04]. Generally speaking, FIUs collect data like threshold transaction reports, international fund transfer instructions, and suspicious matter/activity reports from Reporting Entities (REs), which include banks, money remitters, casinos, law firms, real-estate companies, and financial companies. (In a country the size of Australia, there are about 10,000 REs in the financial system.) FIUs may also get data about entities of interest from partner agencies like law-enforcement agencies and partner FIUs in other countries. Using all these collected data, analysts in FIUs produce intelligence products that either report on specific threats that can be actioned by law enforcement, or describe macro threats and mitigation strategies that financial institutions should adopt in their internal controls. Figure 1 shows the general information flow from reporting entities to FIUs and, from there, to law-enforcement agencies.
Figure 1: A reporting entity is a company that provides financial services that are regulated under AML/CTF laws; KYC = Know Your Customer; ECDD = Enhanced Customer Due Diligence; TTRs = Threshold Transaction Reports; IFTIs = International Fund Transfer Instructions; SMRs = Suspicious Matter Reports; API = Application Programming Interface.
There are safeguards and risk monitoring all over the ecosystem, inside each reporting entity and within the FIU, and there is a reasonable mix of automation and human involvement in those risk-monitoring and risk-reduction processes. The parts of the diagram that are painted in brown are areas of the financial intelligence ecosystem where advanced analytics like predictive and prescriptive modelling can play an important role, mostly because the data volume that need to be processed at those places are enormous and the relationships that need to be mastered and managed, including cause and impact relationships, are often complex.
Generally speaking, all the data received by an FIU are pushed through an entity resolution step to identify and merge records that are highly likely to refer to the same entity [ZNCC18]. This is usually a highly non-trivial computational problem because of the sheer number of parties involved in financial transactions (billions for a typical FIU), to arrive at a table like Figure 2, where the rows are entities (either individuals, companies, or groups of related individuals/companies) and the table show their behaviour in the financial system as seen through the lens of different reporting entities. In some cases, we also have explicit risk indicators for these entities. Given such a table, an FIU equipped with data science expertise and big data analytics platforms can use profiling and statistical machine learning techniques, both supervised and unsupervised at different levels of granularity from individual transactions to single entities and networks of entities, to uncover risks in the financial system and act on them. Examples of such work include [SWC+16, SWZ+17, SZC+17].
Figure 2: The entity-resolved data and risk indicators available to a typical FIU.
A key thing to understand about financial intelligence is that financial crimes cannot really be detected from financial transaction data alone. Here, in a nutshell, are the sort of things we can infer about anomalies in financial transactions [Mad11, RT04]:
  1. 1.
    the source of fund cannot be explained;
  2. 2.
    there are rapid movements of funds through different channels, which may indicate intention to hide the money trail;
  3. 3.
    the financial transaction behaviour is inconsistent with a customer’s expected business profile.
These sorts of activities are not illegal on their own, of course. It is only when these activities accompany a predicate crime — tax evasion, welfare fraud, drug trafficking, etc — that criminality can be established and treated. For that reason, one needs law-enforcement presence throughout the ecosystem to supply leads and the risk contexts for the whole system to work well. In particular, pure discovery work in financial transaction data is quite difficult but commercial entities of interest databases like LSEG World-Check and external datasets like company registration data can be useful in adding context to financial data.
That is the current state. The overall system is far from perfect but, by and large, the regime works reasonably well in deterring, detecting, and disrupting financial crimes. The main limitation is in the links that are not there in the ecosystem. First of all, a lot of financial transaction data, including domestic interbank transfers regardless of amount, are not reported by default to an FIU. Secondly, there are literally tens of thousands of reporting entities in a mature economy like Australia and these reporting entities are largely limited in what intelligence and data they can share with each other because of legal impediments around sharing SMRs between reporting entities to minimise the risk of (accidental) tipping off. There is also not a lot of feedback loops built into the system, including that between reporting entities and an FIU, and that between an FIU and law-enforcement agencies.
The important limitations in the current state can thus be mostly traced to the following two issues:
  • Complex financial crimes typically involve transactions that involve multiple financial institutions in multiple geographies, going through multiple payment channels, some of which are unregulated.
  • FIUs and reporting entities are all working with data silos and no one has a complete view of activities across the financial system.
To appreciate the extent of the problem, one only needs to look at the difficulty of trying to answer a simple question like finding all entities that exhibit a certain cross-border transaction pattern like that shown in Figure 3.
Figure 3: A simple hard problem
The question, and many more complicated others we would like to ask, simply cannot be answered today because the data needed to answer them sit in different organisations in different countries, governed by different legislation.
2.2 Privacy Technologies As A Way Forward
So what can we do? Figure 4 shows a possible way forward using distributed confidential computing technologies, assuming FIUs and partner agencies are willing to work together but without needing to make wholesale changes to AML/CTF legislations everywhere.
Figure 4: Privacy-enhanced version of Figure 2. LEAs = Law Enforcement Agencies; PAs = Partner Agencies; EOI = Entities of Interest
The obvious obstacle to different FIUs and partner agencies sharing large volumes of data with each other is privacy concerns. In the proposed scheme, instead of entity resolution, we have all the participating agencies contributing data that are first matched using privacy-preserving record linkage (PPRL) algorithms [CRS20]. In a PPRL algorithm, the records in different databases that represent or belong to the same entity are identified in a privacy preserving manner without revealing any plaintext values to the participating parties. The matched data are then encrypted using a homomorphic encryption scheme [Hal17] (as described in Section 3.2), which allow mathematical operations to be performed on encrypted data, to arrive at a table similar to what we showed earlier, with two distinctions:
  • The values in the table are all encrypted and the columns of the table are stored in a distributed way across different databases in different agencies.
  • The point of using homomorphic encryption schemes is that it is still possible to run a range of risk-modelling algorithms over the encrypted data, and once we get to a small set of entities of interest, decrypting those entities’ data can be done without raising any privacy concerns for the general public.
Having the data distributed in different data sources does complicate the running of risk-modelling algorithms somewhat but the added security and scalability make it a worthwhile trade-off to make.
There are basically three fundamental computational problems to solve, in increasing order of difficulty:
  1. 1.
    How can two FIUs, each of which has a list of suspicious entities, share only the common suspicious entities with each other, without sharing their entire list with each other? More generally, how can such sharing be done between multiple FIUs?
  2. 2.
    Given a criminal typology that can be described as a graph pattern, where nodes denote entities and edges denote relationships, how can we detect all entities that satisfy the criminal typology when the underlying graph span multiple financial institutions?
  3. 3.
    How can we perform model learning, either supervised or unsupervised, over a set of entities whose financial data and behaviour are spread across multiple financial institutions?
In the rest of this paper, we will look at how aspects of that future state can be realised through different privacy technologies.
3 Privacy Technologies
The purpose of this section is to briefly introduce the most common privacy technologies that are applicable to financial intelligence. These techniques are based on different mathematical foundations to serve different data privacy and security purposes. Secure multiparty computation, through the use of secret sharing schemes, permits multiple parties to jointly compute a function without divulging each party’s secret information to the others. Homomorphic encryption tackles this issue from a different perspective, by encrypting sensitive data in a way that allows arithmetic operations to be performed directly on the ciphertexts. Some privacy issues could arise through reverse-engineering the algorithm or query outcomes [DN10, KM11]. To address this, the differential privacy framework can be used to provide individuals with guaranteed plausible deniability by adding suitably calibrated noise to query outcomes. We will also describe federated learning, which is needed for multiple parties to jointly learn a machine learning model from distributed data. We will start the section with hashing, a foundational technique for many privacy technologies.
3.1 Hashing Techniques
Hashing has a long history in Computer Science. Hashing was first introduced in 1953 to speed up the searching for information in storage systems [Kno75]. The general idea of hashing, or more specifically a hash function, is to map values that are of arbitrary sizes, such as textual strings that can have different lengths, to fixed sized values such as integer numbers that are in a specific range. A hash function therefore converts input values into hash values or hash codes. These hash values are then used as indexes or pointers into a hash table of fixed size. According to [CZ17], a good hash function needs to: (1) map values to hash values as uniformly as possible across the space of possible hash values, (2) the same input value will always result in the same hash value, (3) the mapping process should be fast, and (4) the size of the hash table for a given set of values can be specified by setting the range of integer hash values.
Hashing has been and continues to be a valuable security mechanism for making data unreadable to the human eye, preventing its interception by malicious individuals, and providing a way to validate its integrity. Next we explain few variations of hashing that are useful in the financial intelligence domain.
Keyed Cryptographic Hashing
A keyed hash function (also known as a hash message authentication code, or HMAC) is an algorithm that uses a cryptographic key and a cryptographic hash function to produce a message authentication code that is keyed and hashed [BCK96]. In general HMAC is used to verify the integrity and authenticity of data by computing a fixed-size hash value based on a secret key and input data.
Unlike traditional cryptographic hash functions that generate a hash value solely based on the input data, keyed hashing incorporates a secret key or salt value into the hashing process, making it resistant to tampering and unauthorised modification. The key serves as a shared secret between the sender and receiver, ensuring that only parties possessing the key can compute or verify the hash value. This provides a strong level of security against malicious tampering or forgery attempts, as any alteration to the input data or hash value would require knowledge of the secret key to generate a valid HMAC. Keyed cryptographic hashing is widely used in various security applications, including message authentication, data integrity verification, and secure communication protocols [KSN+23].
Bloom Filters
A Bloom filter [Blo70] is a space-efficient probabilistic data structure used to test whether an element is a member of a set. It is implemented as a bit array $BF[i]$, where $i\in[1,n]$, consisting of $n$ bits. The Bloom filter uses $k$ independent hash functions $H_{1}(x),\cdots,H_{k}(x)$, each of which maps an element $x$ in a set $S$ of $m$ elements to one of the $n$ bits within the bit array $BF$. To determine if an element $x$ is present in the Bloom filter, one needs to check if all $k$ hash functions map $x$ to bits that are set to 1 in the Bloom filter. Because of hash collisions, false positives can occur in a Bloom filter, but false negatives are impossible.
Figure 5: The Dice similarity calculation between the names ‘peter’ and ‘pete’, converted into bigrams ($q=2$) and then encoded into two Bloom filters $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ of length $l=12$ bits using $k=2$ hash functions. The bit “$1$” shown in italics at position $6$ in $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ is a hash collision, because both ‘pe’ and ‘te’ are hashed to this position.
In the privacy-preserving record linkage domain, Bloom filters are commonly used for matching string values or computing similarities between string pairs in a privacy-preserving manner [SBR09, CRS20]. When matching string values, each string is first converted into a set of q-grams, where each q-gram is typically of length 2 or 3. These q-grams are then encoded into a Bloom filter using $k$ hash functions. To compute similarities between Bloom filters, a set-based similarity function such as Jaccard or Dice can be used. Recent literature has also shown that Bloom filters can be used to match both numerical and textual values in databases [WVKR23]. Figure 5 illustrates an example of using a Bloom filter for data matching. The Privacy-Preserving Record Linkage step in Figure 4 can be solved using variations of Bloom Filters. The hash functions in such Bloom Filters are usually keyed cryptographic hashing functions to protect against dictionary attacks.
3.2 Homomorphic Encryption
Homomorphic encryption (HE) is a class of encryption schemes that allows certain arithmetic operations to be performed directly on the ciphertext [AAUC18]. HE schemes are usually public-key cryptosystems. Let $m_{1}$ and $m_{2}$ be two plaintexts, and let $pk$ and $sk$ be the public key and secret key for the encryption Enc() and decryption Dec() functions, respectively. The correctness of a HE scheme is expressed by the following property. For an operation $\diamond$ in the plaintext space, there is a corresponding operation $\bullet$ in the ciphertext space such that
$\text{Dec}(sk,\text{Enc}(pk,m_{1})\bullet\text{Enc}(pk,m_{2}))=m_{1}\diamond m_{2},$ (1)
where the different symbols indicate that the arithmetic operations, either addition or multiplication, in the plaintext and ciphertext spaces may differ. The term “homomorphic” reflects that the decryption function acts like a homomorphism, preserving the structure of the mathematical object.
Depending on the types of arithmetic operations permitted and the number of operations it can handle before decryption fails, an HE scheme is classified into one of the following types, where the algorithmic complexity of the operations we want to perform on data is measured in terms of the depth of circuits and the arithmetic gates supported.
  • Partially HE (PHE) - Schemes that can evaluate circuits containing only one type of arithmetic gate, that is, either addition or multiplication, for unbounded circuit depth.
  • Somewhat HE (SHE) - Schemes that can evaluate a subset of circuits containing both addition and multiplication gates, where the complexity grows as the circuit depth increases.
  • Fully HE (FHE) - Schemes that can evaluate all circuits containing both addition and multiplication gates for unbounded circuit depth.
The privacy-preserving record linkage and model scoring steps in Figure 4 can be done using PHE, but full-scale risk modelling using both supervised and unsupervised machine learning algorithms would usually require SHE or even FHE. FHE was made possible only in recent years due to the breakthrough described in [Gen09]. Gentry’s work is technically based on lattice-based cryptography, an active research area also seen as a strong candidate for quantum-resistant encryption [A+22]. In addition to the standard HE components, Gentry’s scheme introduces two key concepts: squashing and bootstrapping. Simply put, squashing reduces the complexity of decryption by providing a hint to the evaluation process (which introduces additional security assumptions), while bootstrapping allows the scheme to evaluate its own decryption along with an arithmetic operation. When combined, these techniques enable the scheme to perform unbounded arithmetic operations with controlled decryption errors.
It is worth noting that HE schemes are inherently probabilistic, meaning that the same plaintext can produce a different ciphertext each time it is encrypted. This is to ensure a security notion called semantic security. For a more thorough introduction to HE and its mathematical and computer science foundations, the reader is referred to [Hal17, Pei16, LNP22].
To conclude this section, we provide an example of a partially homomorphic encryption (PHE) scheme: the multiplicative HE scheme known as ElGamal Encryption [ElG85]. This scheme is based on the Diffie-Hellman key exchange protocol [DH76] to securely exchange a masked secret key between Alice and Bob for encryption and decryption.
Public key ElGamal Encryption. Bob is the sender, Alice is the receiver. Alice chooses a cyclic group $G$ with order $q$ and a generator $g$.
Private key: Alice’s secret key is a random integer $x\leftarrow\{1,\dots,q-1\}$. Public key: Alice computes $h=g^{x}\in G$ and publishes the public key $(G,q,g,h)$. Encryption: To encrypt a message $m\in G$, Bob randomly selects $y\leftarrow\{1,\dots,p-1\}$, computes the shared secret $s=h^{y}\in G$, and sends the ciphertext $(c_{1},c_{2})$, where $c_{1}=g^{y}$ and $c_{2}=m\cdot s$. Decryption: To decrypt, Alice computes the shared secret $s=c_{1}^{x}$ and its inverse $s^{-1}\in G$, then recovers the message $m=c_{2}\cdot s^{-1}$.
Since each term in the ElGamal ciphertext can be multiplied separately to produce a multiplication ciphertext, and Alice can recover the correct multiplied message using the above decryption process, it implies that ElGamal is a multiplicative HE scheme. Specifically, if $C=(c_{1},c_{2})\leftarrow\text{Enc}(m)$ and $\tilde{C}=(\tilde{c}_{1},\tilde{c}_{2})\leftarrow\text{Enc}(\tilde{m})$, then decryption of the product $C\cdot\tilde{C}$ yields $m\cdot\tilde{m}$, i.e., Dec$(C\cdot\tilde{C})=m\cdot\tilde{m}$. We will see some applications of the additive homomorphic form of ElGamal encryption in Section 4.2.
3.3 Secure Multiparty Computation
Secure multiparty computation (MPC) is a concept that drives the development of protocols designed to securely and accurately perform computations across multiple parties. Research in MPC originated with Yao’s so-called Millionaires’ problem [Yao82], in which two parties seek to know who is richer but without having to disclose their net wealth to each other or a trusted third party. There are many solutions to the Millionaires’ problem involving different cryptographic primitives. Here we briefly describe an accessible solution from [LT05]. The algorithm first converts two numbers in binary representation $x=x_{n}x_{n-1}\ldots x_{1}$ and $y=y_{n}y_{n-1}\ldots y_{1}$ we want to privately compare to the following two sets
$S_{x}=\{x_{n}x_{n-1}\ldots x_{i+1}1\,|\,x_{i}=0,1\leq i\leq n\}$
$S_{y}=\{y_{n}y_{n-1}\ldots y_{i}\,|\,y_{i}=1,1\leq i\leq n\}$
and then securely compute the cardinality $|S_{x}\cap S_{y}|$. The correctness of the procedure comes from the fact that $x>y$ if and only if $S_{x}\cap S_{y}$ is non-empty. There are various private set intersection cardinality (PSI-CA) protocols for computing $|S_{x}\cap S_{y}|$, utilising different techniques such as homomorphic encryption, hashing, Bloom filters, and oblivious transfer. Here, we describe a PSI-CA protocol that employs hashing and commutative encryption from [AES03].
Informally, an encryption function $f$ is commutative if the order of encryption does not matter for arbitrary encryption keys $e$ and $e^{\prime}$, i.e., $f_{e}\circ f_{e^{\prime}}=f_{e^{\prime}}\circ f_{e}$ where $\circ$ is function composition. An example given in [AES03] is the function
$f_{e}(x)=x^{e}\mod p,$
where the powers commute, yielding
$(x^{e}\bmod p)^{e^{\prime}}\bmod p=x^{ee^{\prime}}\bmod p=(x^{e^{\prime}}\bmod p)^{e}\bmod p.$
The security of this commutative encryption is based on the Decisional Diffie-Hellman hypothesis [Bon98]. Armed with a commutative encryption function, $|S_{x}\cap S_{y}|$ can be computed via double encryption as follows:
$|S_{x}\cap S_{y}|=|f_{k_{y}}(f_{k_{x}}(S_{x}))\cap f_{k_{x}}(f_{k_{y}}(S_{y}))|,$
where $k_{x}$ and $k_{y}$ are the encryption keys for the two parties holding the value $x$ and $y$, respectively.
The above secure number comparison protocol provides a glimpse into the rich literature on MPC [Gol98, CD15, EKR+18]. Interestingly, what makes a distributed computation task an MPC protocol cannot be straightforwardly defined in a mathematical way. The well-accepted way of defining MPC’s security is by simulating an ideal world and a real world, known as the ideal/real simulation paradigm [Lin20]. In the simulation, the two worlds are identical except that there is a trusted third party in the ideal world but not in the real world. To examine the security of a protocol under adversarial attacks, one would compare the outcomes of the protocol in the two worlds to determine if the adversary in the real world is able to gather any additional non-disclosed information, should the same attacks also be successful in the ideal world.
With the ideal/real simulation paradigm in place, the following are important properties of an MPC protocol. Privacy guarantees that an adversary learns nothing about an honest party’s input beyond what is prescribed. Correctness ensures that the protocol always produces the correct result despite being attacked by adversaries. Guaranteed output stops the adversary from halting the protocol hence stopping others from receiving an output. Lastly fairness states that an adversary receives the protocol output if and only if an honest party receives the protocol output.
There are two primary adversarial models: semi-honest and malicious adversaries. In the semi-honest model, an adversary correctly follows the protocol specification, but seeks opportunities to learn the protected privacy from information obtained during the computation process. In contrast, the malicious adversarial behaviour can arbitrarily deviate from the protocol, including aborting the protocol or changing its inputs and outputs.
3.4 Differential Privacy
Differential privacy (DP) is a technique that aims to enable the analysis of sensitive data while protecting the privacy of individuals whose data is included in the analysis [DR14]. In general, DP is achieved by adding noise in a controlled manner during the analysis process, ensuring that statistical queries do not reveal information about specific individuals. This allows organisations to derive valuable insights from their data while minimising the risk of unauthorised disclosure or misuse of sensitive personal data.
Definition 3.4.1.
A randomised algorithm $\mathcal{A}$ is $\epsilon$-differentially private if for all neighbouring databases $\mathbf{D}$ and $\mathbf{D^{\prime}}$ that differ in at most one record, and for all sets $\mathcal{O}$ outputs, we have
$Pr[\mathcal{A}(\mathbf{D})\in\mathcal{O}]\leq\exp(\epsilon)\cdot Pr[\mathcal{A}(\mathbf{D^{\prime}})\in\mathcal{O}],$
where $Pr[\cdot]$ denotes the probability of an event.
The parameter $\epsilon$ is known as the privacy budget in DP. A smaller value of $\epsilon$ indicates a stronger privacy guarantee, as it reflects a stricter bound on how much the query results can change in response to the inclusion or exclusion of a single record. The Bayesian semantics of differential privacy can be found in [KS14, KM14, YZN24]; in short, DP guarantees that the odds ratio of an attacker’s prior and posterior beliefs (before and after seeing the query output) in whether a specific record is part of the input database is bounded, and that confers every individual with plausible deniability on whether the person’s data is in the dataset.
An example scenario that illustrates Definition 3.4.1 is differencing attacks. Consider a medical database containing records of individuals who participated in a disease screening. An analyst with access only to counting queries can query the database under different conditions to select two distinct subsets of the data. If the difference between these subsets is designed to isolate a specific individual, the analyst could potentially infer whether or not this individual has the disease by examining the difference between the two query results. Differential privacy prevents this by injecting suitably calibrated noise into the results of these queries, thus preventing the analyst from reliably determining the presence or absence of any particular individual based on the differences in query results.
A fundamental DP technique is the Laplace Mechanism [DMNS06, DR14]. Given a real-valued query function $f:\mathbb{N}^{|\chi|}\rightarrow\mathbb{R}^{k}$, the Laplace Mechanism directly adds Laplace noise to the query outputs as
$M_{L}(\mathbf{D},f,\epsilon)=f(\mathbf{D})+(e_{1},\dots,e_{k}),\text{ where }e_{i}\leftarrow\text{Lap}\left(\frac{\Delta f}{\epsilon}\right),$
where Lap() is the Laplace distribution, and
$\Delta f=\max_{\mathbf{D}\sim\mathbf{D}^{\prime}\in\mathbb{N}^{|\chi|}}{||f(\mathbf{D})-f(\mathbf{D}^{\prime})||_{1}}$
is a measure of the sensitivity of the query on two neighbouring databases $\mathbf{D}$ and $\mathbf{D}^{\prime}$. The independent and identically distributed noise is randomly sampled from the Laplace distribution with mean 0 and scale $\Delta f/\epsilon$. The noise’s scale is proportional to the query sensitivity (when $\epsilon$ is fixed) and inversely proportional to the noise budget $\epsilon$ (when $\Delta f$ is fixed). This makes good sense, since a less sensitive query needs less noise to hide its output variation. A smaller $\epsilon$ indicates stronger privacy guarantee hence more noise needs to be added to the outputs.
In the financial domain, where large volumes of sensitive data are collected and analysed for various purposes such as risk assessment, fraud detection, customer segmentation, and personalised marketing, DP serves as a valuable tool for balancing the need for data-driven insights with privacy protection. By incorporating DP into data analytics workflows, financial institutions can adhere to stringent regulatory requirements, such as GDPR and Basel III [^5] See https://gdpr-info.eu and https://www.bis.org/bcbs/basel3.htm, while still leveraging the richness of their datasets for business intelligence and decision-making [Nis21]. Essentially all financial data are graph data and a survey of DP algorithms for graph data can be found in [LPR+23].
3.5 Federated Learning
Federated learning (FL) [MMR+17] is a new learning paradigm that aims to build a joint machine learning model based on data distributed across multiple sites or owned by different participants. During model training, information such as gradients is exchanged between participants, but the raw data are not shared. The intention is that the exchanged information does not reveal any private data belonging to any party.
Figure 6: A federated learning architecture. In order to guarantee the privacy of the data, federated learning only permits all the data owners to exchange the model gradient ($w$) with the aggregation server. During this process, each data owner trains its own model with the local data, then uploads the local model to the aggregation server. After aggregating all the received models, the server returns the new global model updates ($W$) to each data owner.
Figure 6 illustrates a typical FL process. Initially, each data owner receives a generic global model from the aggregation server. After receiving the initial model, each data owner conducts local training on this model with their own data and then uploads the related gradient information (local model updates) to the aggregation server. The aggregation server averages the updates received from data owners into the global model, then updates the global model to replace each user’s local model. This process repeats until the global model achieves the desired performance or the training reaches the maximum iteration number.
FL can be categorised into three types based on the distribution of data: horizontal FL, vertical FL, and federated transfer learning [LSTS20]. Horizontal FL occurs when the user features of the datasets overlap significantly, but the users themselves overlap minimally. Vertical FL occurs when the user features of the datasets overlap minimally, but the users themselves overlap significantly. In cases where both users and user features of the datasets overlap rarely, transfer learning can be used to address the lack of training data [LKX+20].
Privacy protection is a major concern in FL [ARPS20, NDP+21]. Although FL protects private data on each device by exchanging only model gradients with the aggregation server, the model communication during the entire training process can still potentially leak sensitive information to third parties, for example, through reverse engineering of models [ZLD22]. Two techniques can be employed to prevent private data leakage during FL model transmission:
  • Secure aggregation techniques ensure that raw data or local updates remain private during the aggregation process. Methods such as MPC or HE (described in Sections 3.3 and 3.2) can be used to aggregate updates without revealing individual contributions.
  • DP methods add noise to local updates to protect individual privacy. This noise ensures that the contribution of any single device is indistinguishable from the contributions of others. Aggregating noisy updates provides privacy guarantees while still allowing the central model to learn from the collective data.
In federated learning, model scoring and model learning are two distinct phases of the process. Model scoring refers to the evaluation or assessment of the performance of a machine learning model using a set of evaluation metrics. Common evaluation metrics for model scoring include accuracy, precision, recall, F1-score, and area under the ROC curve (AUC). Both federated model learning and federated model scoring are important steps in the future-state financial intelligence proposed in Figure 4.
4 Privacy-Preserving Information Sharing
In this section, we describe several real-world financial intelligence scenarios wherein information-sharing between different parties is crucial to support effective detection of financial crimes. In each scenario, we discuss the privacy risks during multiparty collaborations and examine the feasibility of applying various privacy-preserving techniques to address these privacy risks.
It is important to recognise that each scenario may entail varying levels of privacy risks. These risks depend on the degree of trust among different parties. The level of trust is basically inversely proportional to the risks of sharing information that have to be shouldered by the collaborating parties, which is in turn a function of the prevailing government legislation (or lack thereof) and applicable penalties in the event of a privacy breach.
The parties involved in conducting financial intelligence analysis today may include both public sectors (domestic or international) and private sectors. Generally speaking, information-sharing between different domestic public sectors involves a high level of trust. In contrast, information sharing between public and private sectors, or among private sectors, often entails a low level of trust. This lack of trust may stem from various factors, including the absence of clear legislation and threats of business competition. For instance, financial institutions tend to refrain from sharing customer transaction data among themselves, despite the potential benefits such information-sharing could offer in enhancing their anti-money laundering controls. Privacy technologies provide the greatest value in low-trust settings that exist either because the parties involved cannot trust each other or a third party for commercial/legal reasons, or when there is ambiguity on whether a public-good sharing of information crosses the threshold into unacceptable privacy intrusion.
4.1 Suspicious Persons List
Problem statement
Suppose an FIU possesses a moderate-sized list of suspicious individuals (or, more generally, commercial entities) obtained, say, from its own investigations and law-enforcement partner agencies and the FIU wishes to acquire additional information about those suspicious individuals, including where they conduct their financial activities and whether they are known to other FIUs. This can be done if the FIU can safely distribute the Suspicious Person List (SPL) to other FIUs and financial institutions to verify the association of individuals on the list with specific financial entities before engaging in further targeted information requests or exchanges with those financial entities.
Privacy risks
If the SPL is shared with the financial institution in a manner that allows for the identification of individuals on the list, there arise the following privacy concerns, especially in the low-trust public-to-private information-sharing setting.
  • (Accidental Tip-off) The financial institution may inadvertently disclose the FIU’s suspicions to an existing customer, prompting the customer to take additional measures to conceal their activities.
  • (Debanking and Displacement) The financial institution may deliberately cease its banking relationship with a suspicious individual before conclusive evidence can be gathered to substantiate or refute the individual’s suspected wrongdoing. This action may be deemed discriminatory if the financial institution’s decision to terminate services is also influenced by personal factors such as the individual’s nationality, race, or name. Further, suspected individuals that are debanked will also move on to other parts of the financial system, including the shadow financial system, and the FIU will lose sight of them.
  • (Need to Know) The FIU can inadvertently disclose the identities of suspicious entities who are not customers of a financial institution, and that would be an unacceptable violation of those individual’s privacy and the need-to-know principle.
Proposed solution
The problem described above of identifying common individuals between the SPL and each participating financial institution’s customer database can be formulated as a private set intersection (PSI) problem. In a two-party computation protocol, the goal of PSI is to identify the common elements that appear in both parties’ sets and nothing else.
We now describe a simple semi-honest PSI protocol from [FNP04] that utilises polynomial evaluation and homomorphic encryption to ensure the protection of unmatched individuals in the SPL. For detailed security and efficiency analyses, readers are encouraged to refer to [FNP04]. For simplicity, it is assumed that each individual can be uniquely identified by an identification (ID) number, which can be derived from a national ID (in some countries) or a combination of personal information. Here are the steps of the protocol.
  1. 1.
    Assuming the FIU’s list of suspicious individuals is $S=\{s_{1},\dots,s_{n}\}$, the FIU first defines the polynomial
    $f(x)=(x-s_{1})(x-s_{2})\cdots(x-s_{n}),$
    whose roots are exactly the elements in the set $S$. The question of whether an arbitrary $s^{\prime}$ is in $S$ can be checked by evaluating $f(s^{\prime})$.
  2. 2.
    The FIU uses polynomial interpolation to find the coefficient representation of $f(x)$
    $f(x)=a_{0}+a_{1}x+\cdots+a_{n}x^{n},$
    which is uniquely defined according to the Fundamental Theorem of Algebra. The FIU then encrypts the coefficients of $f(x)$ using HE and send the encrypted coefficients and the public key to the other parties:
    $\text{Enc}(f)=\{\text{Enc}(a_{0}),\dots,\text{Enc}(a_{n})\}.$
  3. 3.
    Upon receiving the encrypted $f$, each financial institution evaluates the encrypted polynomial $\text{Enc}(f)$ at each element $d\in D$ in its customer database, then multiplies the evaluation result by a random integer $r$ and adds it to $d$. The property of HE guarantees that the result of these operations is equivalent to the following
    $\text{Enc}(rf(d)+d),\forall d\in D.$
    The above formula ensures that if $d$ is a positive match then FIU is able to identify it from the output, and if $d$ is not a positive match then FIU learns nothing about $d$ from the output.
  4. 4.
    Once the matching outputs are shared, FIU decrypts the outputs and find those elements that also appear in its suspicious person list.
By executing this PSI protocol, the FIU and financial institutions collaborate in a confidential manner, enabling the FIU to identify individuals who appear on both the FIU’s SPL and the databases of participating financial institutions. This facilitates the FIU in requesting additional information from appropriate financial institutions to conduct further investigations. Simultaneously, individuals listed on the SPL but absent from a financial institution’s database remain undisclosed to that institution. Similarly, individuals present in a financial institution’s database but not flagged on the SPL remain confidential to the FIU.
Industrial grade solutions
The FIU.net operated jointly by FIUs in the European Union has offered privacy-preserving sharing of suspicious entities for over a decade now, using the Ma3tch algorithm [BM13]. A recent legal review of FIU.net can be found in [Mou20]. In Australia, the Australian Financial Crime Exchange [^6] http://afcx.com.au is a platform for financial institutions to share lists of entities of interest, from a financial crime and cyber crime perspective, with each other and, selectively, with the Australian Taxation Office.
There are now many fast PSI algorithms that can operate at large scale, including [CLR17, RT21, CM20]. A key issue with PSI is that an adversary can use repeated PSI computations to completely recover an organisation’s entire dataset. Proposals like [PLN23] seek to address this issue by introducing DP protection into PSI computations.
4.2 Transaction Flow Tracing and Typology Matching
Problem statement
A key problem in financial crime investigation is tracing the entire digital footprint of “dirty money” within the financial system, allowing investigators to clearly track the movement of funds across different financial institutions, how they are processed and transferred and, crucially, their ultimate usage or destination. We call this problem transaction-flow tracing.
A key task in financial intelligence supported by transaction-flow tracing is the identification of accounts that match certain money-laundering or terrorism-financing (ML/TF) typologies. The NDIS fraud described in [BILC23] serves as an example. Australia’s National Disability Insurance Scheme (NDIS) provides financial support to eligible participants to improve their quality of life. However, continuous fraudulent activities against the NDIS scheme have resulted in several billions of dollars in losses from its intended use [AUS20]. These fraudulent transactions share common patterns, crossing multiple organisations and eventually being transferred overseas. No single institution in the financial system would have all the data required to identify accounts and entities that exhibit such behaviour. We will show in this section how privacy technologies can help tackle this broad class of problems.
Privacy risks
This scenario exemplifies a case of government-to-private and private-to-private information sharing, which are generally low-trust settings. Complex ML/TF typologies can usually be described as a graph pattern, where nodes correspond to accounts/entities and edges correspond to transactions between those nodes and there can be non-trivial conditions attached to the nodes and edges. To enable the identification of entities that match a complex ML/TF typology that span the financial system, we would need accounts and transaction data from multiple financial institutions to be matched at scale before a filtering or sub-graph matching operation is performed. This is obviously highly intrusive to entities who are not involved in financial crimes, so we need a way of performing the large-scale data-matching and graph pattern searching in a privacy-preserving manner.
Proposed solution
The FinTracer solution proposed in [BILC23] is basically a way to perform an efficient and privacy-preserving breadth-first search across a large distributed graph in homomorphically-encrypted space, allowing Financial Intelligence Units like AUSTRAC to work with law-enforcement agencies and financial institutions to start with some known suspicious entities and find their footprint across the financial system
  1. 1.
    without having to disclose the suspicious entities to financial institutions, thus avoiding the risk of accidentally tipping off the entities; and
  2. 2.
    without having to violate the privacy of unrelated entities even though their data need to be processed.
The key observation behind this algorithm is that large distributed graphs can be represented using sparse matrices and graph traversal can be expressed as matrix multiplication operations [Bur21]. This basic graph search algorithm can serve as a building block to detect a large range of financial crime typologies (i.e. graph patterns), thus significantly advancing the state-of-the-art in intelligence agency’s ability to detect complex financial crimes in a responsible, privacy-preserving manner at a large scale.
[…]
Figure 7: A simplified transaction network to demonstrate FinTracer on the NDIS example. Accounts with cross organisation transactions are linked by directed edges. There are four financial institutions A, B, C and D, each having a number of accounts. The account a1 in A directly received NDIS payment.
To see how this works, suppose AUSTRAC is interested in identifying the accounts (from multiple financial institutions) that transferred NDIS funds overseas. These accounts either received funds directly from NDIS or indirectly via a limited number of cross-institution transactions. The FinTracer tag $t_{\leq}$ is implemented to keep track of accounts with at most $k$ cross-institution transactions. These tags are homomorphically encrypted to securely follow NDIS funds in the entire transaction network. The key components of the FinTracer process are the ElGamal encryption scheme [ElG85] over the additive Curve25519 for its performance efficiency [Ber06], and the aggregations of the tag values across financial institutions over multiple iterations of the breadth-first search algorithm.
Figure 7 is a simplified example of the NDIS fraud scenario. In this example, there are four financial institutions. The only account that directly received suspicious NDIS fund is a1 in institution A. All institutions are only aware of the source and destination accounts that are involved in the transactions they process. In its simplest form (Algorithms 1 and 2 [BILC23]), the FinTracer protocol does the following steps:
  • Initialisation: Each financial institution sets the initial FinTracer tag of each of its customer account $x$ to either $\text{Enc}(1)$ or $\text{Enc}(0)$, depending on whether $x$ receive NDIS fund directly. In the example above,
    • $t_{\leq}(a1)=\text{Enc}(1)$ and $t_{\leq}(\{b1,b2,c1,c2,d1\})=\text{Enc}(0)$ (the set notation denotes that the tag is set for element in the set).
    • The other accounts do not have cross-institution transactions, hence are out of considerations by the protocol.
  • Iteration 1: In the first iteration, the initialised tags are passed to the neighbouring institutions following the transaction flow. The receiving institutions then aggregate the tags with the received tags. In the example, the shared partial mappings are
    • A $\rightarrow$ B: $P^{A}_{B}((a1,b1))=\text{Enc}(1),P^{A}_{B}((a1,b2))=\text{Enc}(1)$.
    • B $\rightarrow$ C: $P^{B}_{C}((b1,c1))=\text{Enc}(0),P^{B}_{C}((b2,c2))=\text{Enc}(0)$.
    • C $\rightarrow$ D: $P^{C}_{D}((c1,d1))=\text{Enc}(0),P^{C}_{D}((c2,d1))=\text{Enc}(0)$.
    After receiving these partial mappings, the aggregated tags are
    • $t_{\leq}(\{a1,b1,b2\})=\text{Enc}(1)$ and $t_{\leq}(\{c1,c2,d1\})=\text{Enc}(0)$
    since $t_{\leq}(b1)=t_{\leq}(b1)+P^{A}_{B}((a1,b1))=\text{Enc}(0)+\text{Enc}(1)=\text{Enc}(1)$ and similarly for $t_{\leq}(b2)$. The tag values of $c1$, $c2$, $d1$ remain $\text{Enc}(0)$ since the initial and received tags are $\text{Enc}(0)$.
  • Iteration 2: The partial mappings are updated by institution B yielding
    • A $\rightarrow$ B: $P^{A}_{B}((a1,b1))=\text{Enc}(1),P^{A}_{B}((a1,b2))=\text{Enc}(1)$.
    • B $\rightarrow$ C: $P^{B}_{C}((b1,c1))=\text{Enc}(1),P^{B}_{C}((b2,c2))=\text{Enc}(1)$.
    • C $\rightarrow$ D: $P^{C}_{D}((c1,d1))=\text{Enc}(0),P^{C}_{D}((c2,d1))=\text{Enc}(0)$.
    After receiving these partial mappings, the aggregated tags are
    • $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(2)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(1)$ and $t_{\leq}(d1)=\text{Enc}(0)$.
  • Iteration 3: The partial mappings are updated by B and C yielding
    • A $\rightarrow$ B: $P^{A}_{B}((a1,b1))=\text{Enc}(1),P^{A}_{B}((a1,b2))=\text{Enc}(1)$.
    • B $\rightarrow$ C: $P^{B}_{C}((b1,c1))=\text{Enc}(2),P^{B}_{C}((b2,c2))=\text{Enc}(2)$.
    • C $\rightarrow$ D: $P^{C}_{D}((c1,d1))=\text{Enc}(1),P^{C}_{D}((c2,d1))=\text{Enc}(1)$.
    After receiving these partial mappings, the aggregated tags are
    • $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(3)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(3)$ and $t_{\leq}(d1)=\text{Enc}(2)$.
The FinTracer tags are now sent to the FIU and decrypted. Only those accounts with a non-zero tag have received money that can be traced either directly or indirectly to suspicious NDIS payment.
Industrial grade solutions
An open-source, scalable implementation of FinTracer can be found at http://github.com/AUSTRAC. The implementation includes a C/Cuda library for additive HE with ElGamal over the Ed25519 elliptic curve. A number of algorithms for different graph operations are included in http://github.com/AUSTRAC/ftillite/Documentation/Purgles. The extension of FinTracer, which allows only linear operations because of the use of additive HE, to non-linear set and graph operations to allow the detection of more-or-less arbitrarily complex criminal typologies is described in [BCF23]. Note that differential privacy noise is added in several steps in these algorithms to avoid variations of database-reconstruction attacks [DN03, DMT07].
Finding key players in transaction networks
Once the small number of accounts and entities that matched a certain ML/TF typology is detected using a technology like FinTracer, an FIU like AUSTRAC has the legal power to request much more comprehensive information about those accounts and entities from financial institutions because the request for information or notice to produce information is now targeted and backed by genuine grounds for suspicion. (The remaining small-scale privacy concerns can mostly be addressed using differentially private graph search algorithms like [KRWY16].) With such a dataset, it then makes sense to calculate connected components and risk metrics like PageRank scores and betweenness centrality scores, some of which have been shown to be useful features to enhance the accuracy or confidence of machine learning or rule-based financial intelligence classification models [MCF+17, SvHA+19, vEDvdB+24].
4.3 Collaborative Model Learning
Problem statement
The use of supervised learning and anomaly detection algorithms to learn financial crime models from data has long been a topic of interest within financial intelligence [SvHA+19, vEDvdB+24], but the practical utility of such approaches has been hampered by a lack of quality labelled data in the case of supervised learning algorithms, and high false positive rates in the case of anomaly detection algorithms.
Recent advances in privacy-preserving FL have the potential to alleviate some of these issues by allowing different business units within a financial institution and even multiple financial institutions to jointly learn models from their collective data. In this section, we describe some of the key ideas behind some exploratory work in this area.
Privacy risks
To enable FL, we would need accounts and transaction data from multiple financial institutions to be matched, often probabilistically, at scale before a machine learning algorithm can be run. In the case of supervised learning, the participating parties would also need to share, as labelled data, entities that are either known or suspected to be engaging in financial crime. This is obviously highly intrusive to entities who are not involved in financial crimes, and also causes the kind of tipping-off issues discussed in Section 4.1 for those (suspected to be) involved in financial crimes. All these imply that we need scalable and secure ways to perform large-scale probabilistic data matching and model learning in a privacy-preserving manner. The publication and reuse of machine learning models by multiple organisations also open the door to attacks like membership inference that need to be managed [MPP+21].
Proposed solutions
As discussed, the first step in FL is privacy-preserving data matching. Unlike transaction flow tracing, where the source and target of each transaction are uniquely identifiable for exact matching, there can be inconsistencies in how different financial institutions collect and store individual information, such as names, addresses, driver’s licences, business names, and other details. Consequently, data matching and entity resolution in federated learning – both within and across financial institutions – are often performed probabilistically. A comprehensive description of key privacy-preserving data matching/record linkage algorithms, including those that are built on Bloom Filters as described in Section 3.1, can be found in [CRS20].
We now describe a simplified example of federated learning to illustrate its applicability in financial intelligence. The task is to learn a linear regression risk model using probabilistically matched data from multiple participating parties, while preserving privacy in a distributed manner. The target of the regression is the individual risk score, and the features for the regression are derived from all the data held by the REs; the setup is shown in Figure 8 for three parties.
More formally, given a training set $\{({\bf\it x}_{l},y_{l})\}_{l=1,\dots,m}$ that includes $m$ individuals, where the feature set ${\bf\it x}_{l}={\bf\it x}_{l}^{(1)}\cup\cdots\cup{\bf\it x}_{l}^{(p)}$ is the union of the features from all parties, the task is to fit a linear regression model
${\bf\it Y}={\bf\it X}{\bf\it w}+{\bf\it\epsilon}$
that predicts the target variable ${\bf\it Y}$ as accurate as possible.
Figure 8: Federated learning setup for privacy-preserving risk modelling
The regression training is performed using gradient descent. The weight vector ${\bf\it w}$ is initially assigned random values and then updated using the gradient of the loss function with respect to each weight element $w_{j}$ as follows, until convergence.
$w_{j}:=w_{j}-\alpha\frac{1}{m}\sum_{l=1}^{m}(\langle{\bf\it x}_{l},{\bf\it w}\rangle-y_{l})x_{l,j}$ (2)
The algorithm converges to the optimal solution because the least-squares loss function is convex and the learning rate $\alpha$ controls the convergence rate.
Below are the steps for the privacy-preserving distributed version of the algorithm that uses an additive HE system.
  1. 1.
    The FIU generates a public-private key pair ($pk$, $sk$) and shares the public key $pk$ with the REs.
  2. 2.
    Each RE initialises its portion of the weight coefficients ${\bf\it w}^{(i)}$ randomly and computes a partial prediction for each individual
    ${\bf\it Y}^{(1)}$ $=(\langle{\bf\it x}^{(1)}_{1},{\bf\it w}^{(1)}\rangle,\langle{\bf\it x}^{(1)}_{2},{\bf\it w}^{(1)}\rangle,\ldots,\langle{\bf\it x}^{(1)}_{m},{\bf\it w}^{(1)}\rangle),$
    ${\bf\it Y}^{(2)}$ $=(\langle{\bf\it x}^{(2)}_{1},{\bf\it w}^{(2)}\rangle,\langle{\bf\it x}^{(2)}_{2},{\bf\it w}^{(2)}\rangle,\ldots,\langle{\bf\it x}^{(2)}_{m},{\bf\it w}^{(2)}\rangle),$
    $\vdots$
    ${\bf\it Y}^{(p)}$ $=(\langle{\bf\it x}^{(p)}_{1},{\bf\it w}^{(p)}\rangle,\langle{\bf\it x}^{(p)}_{2},{\bf\it w}^{(p)}\rangle,\ldots,\langle{\bf\it x}^{(p)}_{m},{\bf\it w}^{(p)}\rangle),$
    where ${\bf\it x}^{(i)}_{l}$ is the feature set of individual $l$ held by the $i^{th}$ RE.
  3. 3.
    Each RE homomorphically encrypts its partial prediction element-wise and participates in a round robin communication to aggregate a complete prediction for each individual. The resulting encrypted prediction is then sent to the FIU for subsequent computation.
    $\overline{{\bf\it Y}}=\left(\sum_{i=1}^{p}\text{Enc}(pk,{\bf\it Y}^{(i)}_{1}),\dots,\sum_{i=1}^{p}\text{Enc}(pk,{\bf\it Y}^{(i)}_{m})\right),$
  4. 4.
    The FIU decrypts the complete prediction and computes the difference from the actual values
    ${\bf\it E}=\text{Dec}(sk,\overline{{\bf\it Y}})-{\bf\it Y},$
    and then sends the unencrypted ${\bf\it E}$ back to the REs.
  5. 5.
    Each RE then updates its portion of the weight vector ${\bf\it w}^{(i)}$ independently using Equation 2
    ${\bf\it w}^{(i)}_{j}={\bf\it w}^{(i)}_{j}-\alpha\frac{1}{m}\sum_{l=1}^{m}{\bf\it E}\cdot x^{(i)}_{l,j}$
    where $j$ ranges over the variables held by the RE.
  6. 6.
    Repeat steps 2 to 5 until the model has converged.
Throughout the entire computation, the FIU only sees the predicted values from the REs, from which it cannot recover the data held by the REs. Conversely, the REs only see the encrypted partial predictions from each other and the gradients from the FIU, from which they cannot recover the data held by the FIU or other REs.
In the remainder of this subsection, we provide a survey of some innovative use of FL that holds promise for financial intelligence. In the work of [SZK+22], we see a reasonably compelling demonstration of how FL can significantly enhance the detection capabilities of financial crime models. Through a series of case studies and rigorous experimental evaluations, this research highlights the dual achievement of improved detection performance and robust data privacy. The study underscores how FL can elevate the precision of financial crime detection while safeguarding sensitive information.
Building upon this foundation, [DNM+23] introduces a novel approach that advances secure and collaborative model training across multiple financial institutions. By integrating HE with FL, this approach achieves secure model aggregation and parameter updates. The experimental results and evaluations presented reveal the approach’s effectiveness in preserving data privacy and security, thus fostering a collaborative environment for financial data analysis across institutions.
Further extending this narrative, [HHIL+17] presents an end-to-end scheme designed for linear classifiers, employing additive HE to counteract threats from semi-honest adversaries in vertical FL scenarios. This research delves into the impact of entity resolution on learning performance, utilising a Taylor approximation for the loss and gradient functions. The application of Paillier additive HE within this context enables privacy-preserving computations, showcasing an innovative method for handling sensitive data securely.
Additionally, [BP20] proposes a FL scheme tailored for fraud detection. In this scheme, individual financial organisations enhance their model parameters with DP noise before transmitting them to a central server. This strategy effectively diminishes the server’s ability to recover private information, advancing the cause of privacy in fraud detection efforts.
Though these FL techniques have yet to achieve industrial-grade maturity, they collectively offer a glimpse into the forefront of financial intelligence. Their contributions highlight ongoing advancements and set the stage for future innovations in the field.
Industrial grade solutions
There are several industrial-grade solutions like [ZCN17, ZNCC18] for conducting probabilistic matching of databases with millions to billions of entities and these can be used in conjunction with techniques like Bloom filters and HE to produce somewhat scalable privacy-preserving probabilistic data-matching algorithms. While federated learning is now a mature technology, progress in using federated learning at the industrial scale for financial intelligence applications has been slow because, in most cases, there is not an abundance of labelled data to support the acquisition of robust and trusted machine-learning models with low false-positive rates. With further advances in FL and a general uplift in the maturity of FIUs in adopting privacy technologies, we expect to see more progress in this area in the future.
5 Discussion and Conclusion
Financial intelligence equips law enforcement and financial institutions with sophisticated analytical tools and data-driven methodologies to proactively detect and prevent financial crimes. While financial intelligence plays a crucial role in supporting Anti-Money Laundering (AML) and Counter-Terrorist Financing (CTF) regimes by delivering actionable intelligence and insights, it may present challenges in relation to data privacy regulations such as Australia’s Privacy Act and European Union’s General Data Protection Regulation (GDPR). These challenges arise particularly when there is a discrepancy between the initial purpose for which data was collected by financial institutions and its subsequent use in various financial intelligence applications. While public and private sectors, industry associations and regulatory bodies are actively working to harmonise AML/CTF and data privacy regulations, such changes often involve lengthy and uncertain political approval and implementation processes.
From the practical perspective of being able to do something about complex financial crimes today, we advocate for the integration of privacy-preserving technologies within financial intelligence tools, not just as a mitigation strategy while waiting for legal experts to navigate the complex AML/CTF reform landscape [Bug18], but to inform those AML/CTF reform discussions. This approach offers several benefits. First, privacy-preserving technologies can be swiftly implemented, allowing financial institutions and financial intelligence units to address data privacy concerns while fulfilling AML/CTF obligations. Second, adopting privacy-preserving methods helps institutions build trust with customers and stakeholders by demonstrating a commitment to both data protection and effective financial crime prevention. Third, privacy-preserving solutions offer a flexible framework that can adapt to evolving regulatory requirements, helping institutions stay compliant as laws and regulations are updated.
We hope this paper enhances public understanding of privacy-preserving financial intelligence through the various AML/CTF use cases we explore. The use cases highlight the evolving nature of financial crimes and the necessity for ongoing innovation in privacy technologies to address emerging threats effectively. Having said that, for financial institutions and regulatory bodies, adopting privacy technologies involves careful consideration of both technological and operational aspects. Effective implementation strategies must balance the trade offs between data privacy and the quality and efficiency of analysis. While these technologies address immediate challenges, they are not a substitute for comprehensive regulatory reforms. Continuous refinement of regulations is essential to align financial crime prevention efforts with data privacy goals.
Acknowledgements
We would like to thank the many former colleagues in AUSTRAC and financial crime units in major Australian financial institutions who spent many hours teaching us the intricacies of the financial intelligence business. We are also grateful to many of them for continuing to play an important role to keep our financial system and our country safe.

### v2

@@
 Abstract
-Financial crimes like terrorism financing and money laundering can have real impacts on society, including the abuse and mismanagement of public funds, increase in societal problems such as drug trafficking and illicit gambling with attendant economic costs, and loss of innocent lives in the case of terrorism activities.
-Complex financial crimes can be hard to detect primarily because data related to different pieces of the overall puzzle is usually distributed across a network of financial institutions, regulators, and law-enforcement agencies and they cannot be easily shared due to privacy constraints.
-Recent advances in Privacy-Preserving Data Matching and Machine Learning provide an opportunity for regulators and the financial industry to come together to solve the risk-discovery problem with technology.
-This paper provides a survey of the financial intelligence landscape and where opportunities lie for privacy technologies to improve the state-of-the-art in financial-crime detection.
+Financial crimes like money laundering and terrorism financing can have significant impacts on society, including loss of trust in the integrity of the financial system, misuse and mismanagement of public funds, increase in societal problems like drug trafficking and illicit gambling, and loss of innocent lives due to terrorism activities.
+Effective detection of complex financial crimes remains a formidable challenge for regulators and financial institutions because the critical data needed to establish patterns and criminality are often dispersed across multiple organisations and cannot be linked due to privacy constraints around large-scale data matching.
+Recent advances in privacy and confidential computing technologies, which enable private and secure data analysis across organisations, offer a promising opportunity for regulators and the financial industry to come together to enhance their collaborative risk detection while maintaining privacy standards.
+This paper, through a survey of the financial intelligence ecosystem, seeks to identify opportunities for the utilisation of privacy technologies to improve the state-of-the-art in financial-crime detection.
+keywords:
+Financial intelligence , Anti-money laundering , Privacy technology
+[…]
+[inst1]organization=School of Information Technology, Deakin University, state=VIC, country=Australia
+[…]
+[inst2]organization=Data61, CSIRO, state=ACT, country=Australia
+[…]
+[inst3]organization=School of Computing, Australian National University, state=ACT, country=Australia
 1 Introduction
-Our economy depends on a well functioning banking and financial system that has the trust of the general population.
-However in recent years, the reputation of some major banks has been severely damaged as criminal elements have exploited systemic weaknesses to launder huge sums of money, including transfers to terrorist organisations [Lyn22, LF23].
-Besides reputational damage to financial institutions, complex financial crimes like terrorism financing and money laundering can have real direct impacts on society too, including
-  • abuse and mismanagement of public funds;
-  • increase in societal problems like drug trafficking, illicit trade and gambling, people smuggling, with attendant economic costs; and
-  • loss of innocent lives in the case of terrorism activities.
-Complex financial crimes can be hard to detect primarily because data related to different pieces of the overall puzzle are usually distributed across a network of financial institutions, regulators, and law-enforcement agencies [LCSH19].
-The problem is also rapidly increasing in complexity because of new emerging value-transfer platforms like crypto-currencies and real-time settlement systems.
-In early 2016, AUSTRAC, Australia’s Financial Intelligence Agency, set up the Fintel Alliance [^1] https://www.austrac.gov.au/partners/fintel-alliance, an organisation that includes AUSTRAC and the major financial institutions and law-enforcement agencies, to improve public-private collaboration in combating financial crimes.
-Since then, financial intelligence agencies around the world have started building public-private data-sharing platforms to share and disrupt criminal entities engaging in financial crimes, including the Joint Money Laundering Intelligence Task Force (JMLIT) [^2] https://www.nationalcrimeagency.gov.uk/what-we-do/national-economic-crime-centre in UK, and the Financial Crime Enforcement Network Exchange [^3] https://www.fincen.gov/resources/financial-crime-enforcement-network-exchange in the United States.
-These public-private platforms are highly effective against known threats.
-However, their discovery functions to find as-yet unknown risks in the financial system are still not as strong as they can be, primarily because of privacy constraints around large-scale data matching and analytics between organisations.
-Recent advances in Privacy-Preserving Data Matching and Machine Learning provide an opportunity for regulators and the financial industry to come together to solve the risk-discovery problem with technology.
-This paper provides a survey of the financial intelligence landscape and where opportunities lie for privacy technologies to improve the state-of-the-art in financial-crime detection.
-Our paper complements the material covered in [Max20], which explored global case studies of public and private sectors employing cryptographic-based privacy technologies to foster collaborative financial intelligence.
-[^4] https://www.future-fis.com/the-pet-project.html
+A robust and trustworthy banking and financial system is essential for the stability and proper functioning of any mature economy.
+However, in recent years, the reputation and integrity of several major financial institutions have been seriously undermined as organised criminal networks have exploited loopholes and systemic vulnerabilities to launder vast sums of money, some of which have even been linked to terrorist financing [39, 37].
+Beyond the reputational harm inflicted on financial institutions, sophisticated financial crimes like money laundering and terrorism financing have profound societal consequences, including:
+  • 1.
+    mismanagement and misuse of public funds;
+  • 2.
+    financial losses for institutions due to increased compliance costs and regulatory penalties;
+  • 3.
+    escalation of societal issues such as drug trafficking, human smuggling, and illegal gambling, which carry significant economic burdens;
+  • 4.
+    the tragic loss of innocent lives in the case of terrorist activities.
+Globally important cases like Binance’s \$4 billion fine [^1] Binance’s \$4 billion fine for failing to maintain an effective anti-money-laundering (AML) program and other violations against the US Bank Secrecy Act, UBS’s \$4.2 billion fine by Swiss, US and French regulators for facilitating tax evasion and inadequate AML controls, and Goldman Sachs’s \$2.9 billion fine [^2] Goldman Sachs’s \$2.9 billion fine for the 1MDB money-laundering and bribery violations, have certainly brought these issues into the public consciousness and discourse.
+In Australia, the government’s successful court proceedings against the Commonwealth Bank of Australia, Westpac Banking Corporation, and Crown Casinos for serious breaches of the AML/CTF Act, and their ongoing legal actions against the operators of the Star Casinos and online betting sites like Ladbrokes and Neds highlight the continued challenges faced by regulators and financial intelligence agencies in combating financial crimes.
+[^3] Lists of enforcement actions taken
+In response to these challenges, AUSTRAC, Australia’s Financial Intelligence Agency, established the Fintel Alliance [^4] Fintel Alliance in 2016 to foster collaboration between government agencies, financial institutions, and law enforcement.
+A number of similar initiatives have also emerged globally, including the UK’s Joint Money Laundering Intelligence Task Force [^5] UK’s Joint Money Laundering Intelligence Task Force and the U.S.
+Financial Crime Enforcement Network Exchange [^6] U.S.
+Financial Crime Enforcement Network Exchange.
+These partnerships have proven effective in addressing known threats, but they still face formidable challenges in detecting complex financial crimes because the critical data needed to establish patterns and criminality are often dispersed across multiple organisations and cannot be linked due to privacy constraints around large-scale data matching [36].
+The problem is further compounded by the growing complexity of emerging value-transfer mechanisms such as cryptocurrencies and real-time payment systems.
+Recent innovations in Privacy-Preserving Data Matching and Machine Learning, which enable private and secure data analysis across organisations, offer a promising opportunity for regulators and the financial industry to come together to enhance their collaborative risk detection while maintaining privacy standards.
+This paper, through a survey of the financial intelligence ecosystem, seeks to identify opportunities for the utilisation of privacy technologies to improve the state-of-the-art in financial-crime detection.
+Our paper complements the material covered in [41], which explored global case studies of public and private sectors employing privacy technologies to foster collaborative financial intelligence.
 2 Financial Intelligence
 Financial intelligence is a highly specialised field that is not well understood by many.
-We provide an introduction to the financial intelligence ecosystem in this section, including a description of the current state, some existing limitations and opportunities.
+This section provides an introduction to the financial intelligence ecosystem, including a description of the current state, some existing limitations and opportunities.
 2.1 The Financial Intelligence Ecosystem
-Financial Intelligence Units (FIUs) around the world are governed by anti-money laundering and counter-terrorism financing (AML/CTF) regimes that are quite similar in nature [Mad11, Gil04].
-Generally speaking, FIUs collect data like threshold transaction reports, international fund transfer instructions, and suspicious matter/activity reports from Reporting Entities (REs), which include banks, money remitters, casinos, law firms, real-estate companies, and financial companies.
-(In a country the size of Australia, there are about 10,000 REs in the financial system.) FIUs may also get data about entities of interest from partner agencies like law-enforcement agencies and partner FIUs in other countries.
-Using all these collected data, analysts in FIUs produce intelligence products that either report on specific threats that can be actioned by law enforcement, or describe macro threats and mitigation strategies that financial institutions should adopt in their internal controls.
-Figure 1 shows the general information flow from reporting entities to FIUs and, from there, to law-enforcement agencies.
-Figure 1: A reporting entity is a company that provides financial services that are regulated under AML/CTF laws; KYC = Know Your Customer; ECDD = Enhanced Customer Due Diligence; TTRs = Threshold Transaction Reports; IFTIs = International Fund Transfer Instructions; SMRs = Suspicious Matter Reports; API = Application Programming Interface.
-There are safeguards and risk monitoring all over the ecosystem, inside each reporting entity and within the FIU, and there is a reasonable mix of automation and human involvement in those risk-monitoring and risk-reduction processes.
-The parts of the diagram that are painted in brown are areas of the financial intelligence ecosystem where advanced analytics like predictive and prescriptive modelling can play an important role, mostly because the data volume that need to be processed at those places are enormous and the relationships that need to be mastered and managed, including cause and impact relationships, are often complex.
-Generally speaking, all the data received by an FIU are pushed through an entity resolution step to identify and merge records that are highly likely to refer to the same entity [ZNCC18].
-This is usually a highly non-trivial computational problem because of the sheer number of parties involved in financial transactions (billions for a typical FIU), to arrive at a table like Figure 2, where the rows are entities (either individuals, companies, or groups of related individuals/companies) and the table show their behaviour in the financial system as seen through the lens of different reporting entities.
-In some cases, we also have explicit risk indicators for these entities.
-Given such a table, an FIU equipped with data science expertise and big data analytics platforms can use profiling and statistical machine learning techniques, both supervised and unsupervised at different levels of granularity from individual transactions to single entities and networks of entities, to uncover risks in the financial system and act on them.
-Examples of such work include [SWC+16, SWZ+17, SZC+17].
-Figure 2: The entity-resolved data and risk indicators available to a typical FIU.
-A key thing to understand about financial intelligence is that financial crimes cannot really be detected from financial transaction data alone.
-Here, in a nutshell, are the sort of things we can infer about anomalies in financial transactions [Mad11, RT04]:
+Financial Intelligence Units (FIUs) around the world are governed by anti-money laundering and counter-terrorism financing (AML/CTF) regimes that are quite similar in nature [40, 28].
+Reporting Entities (REs) are companies that provide designated financial services that are regulated under AML/CTF laws and typically include banks, money remitters, digital currency exchanges, casinos, and increasingly law firms and real-estate companies.
+(In a country the size of Australia, there are about 10,000 REs in the financial system.) These REs are required by law to maintain know-your-customer (KYC) and anti-money laundering programs to make sure they can verify their customers’ identities and prevent their designated services from being used for illicit purposes.
+Certain types of customer transaction data, including threshold transaction reports (TTRs) for cash transactions exceeding \$10k and all international fund transfer instructions (IFTIs) regardless of amount, are automatically sent to the relevant FIUs on a regular basis.
+In addition, REs can also submit suspicious matter reports (SMRs) or suspicious activity reports (SARs) to FIUs if they detect suspicious customer activities through their internal controls.
+Adding to all these data from REs, FIUs may also get reports about specific individuals or entities of interest from law-enforcement agencies and partner FIUs in other countries.
+Using all these collected data, analysts in FIUs produce intelligence products that either report on specific threats that can be actioned by law enforcement agents, or describe macro threats and mitigation strategies that financial institutions should adopt in their internal controls.
+Figure 1 shows the information flow from reporting entities to FIUs and, from there, to law-enforcement agencies and the limited feedback loops in the system.
+Figure 1: The figure shows a simplified schematic view of a financial intelligence ecosystem.
+As we can see, there are risk-monitoring and risk-reduction mechanisms built into many different parts of the financial intelligence ecosystem, both within reporting entities in a distributed way and more centrally within the FIU.
+Many of these organisations also have large-scale automated transaction-monitoring systems that are augmented by manual analyst investigations and formal governance oversight.
+The parts of Figure 1 coloured in brown represent areas of the financial intelligence ecosystem where artificial intelligence and machine learning can have a significant impact.
+This is primarily because the sheer quantity of data that must be processed in these locations is immense, and the connections and relationships that need to be distilled and understood, including causal and deliberately hidden linkages, are often complex and intricate.
+Broadly speaking, all data received by an FIU are typically first processed through an entity-resolution step to identify and consolidate records that are highly probable to relate to the same underlying entity [65].
+The aim is to produce a dataset like that shown in Figure 2, where each row represents an entity (either persons, corporations, or clusters of related individuals/companies) and the columns illustrate their activities within the financial ecosystem as observed through the lens of the financial institutions whose services they engage with.
+A small set of such entities would also have explicit risk indicators flagged against them based on, for example, suspicious matter reports, sanctions list and other criminal intelligence data.
+Given a dataset like Figure 2, an FIU equipped with appropriate technical expertise and advanced data-analytics platforms can employ profiling and statistical machine learning methodologies, both supervised and unsupervised, at varying levels of granularity from individual transactions to single entities and networks of entities, to try to uncover potential risks in the financial system and take appropriate action.
+Examples of such work include [55, 56, 57].
+Figure 2: The entity-resolved data and risk indicators available to a typical financial intelligence agency.
+A crucial aspect to understand about financial intelligence is that financial crimes cannot be definitively identified solely from financial transaction data.
+Here, at a high level, are the types of anomalies we can infer from financial transaction data [40, 52]:
   1. 1.
-    the source of fund cannot be explained;
+    the origin of funds cannot be explained or is obscure;
   2. 2.
-    there are rapid movements of funds through different channels, which may indicate intention to hide the money trail;
+    there are rapid transfers of money through various channels, potentially indicating an attempt to obscure the money trail;
   3. 3.
-    the financial transaction behaviour is inconsistent with a customer’s expected business profile.
-These sorts of activities are not illegal on their own, of course.
-It is only when these activities accompany a predicate crime — tax evasion, welfare fraud, drug trafficking, etc — that criminality can be established and treated.
-For that reason, one needs law-enforcement presence throughout the ecosystem to supply leads and the risk contexts for the whole system to work well.
-In particular, pure discovery work in financial transaction data is quite difficult but commercial entities of interest databases like LSEG World-Check and external datasets like company registration data can be useful in adding context to financial data.
+    the financial transaction patterns are inconsistent with a customer’s anticipated personal or business profile.
+It is important to note that these activities, in isolation, are not inherently illegal.
+Criminality can only be established and addressed when these activities are linked to a predicate crime, such as tax evasion, welfare-payment fraud, drug trafficking, or similar crimes.
+To ensure the entire system functions effectively, law enforcement presence is essential throughout the financial intelligence ecosystem to provide leads and supply appropriate risk contexts.
+Specifically, pure discovery work within financial transaction data presents considerable challenges, but commercial databases of entities of interest, such as LSEG World-Check, and external datasets like company registration information can prove valuable in adding context to financial data.
 That is the current state.
-The overall system is far from perfect but, by and large, the regime works reasonably well in deterring, detecting, and disrupting financial crimes.
-The main limitation is in the links that are not there in the ecosystem.
-First of all, a lot of financial transaction data, including domestic interbank transfers regardless of amount, are not reported by default to an FIU.
-Secondly, there are literally tens of thousands of reporting entities in a mature economy like Australia and these reporting entities are largely limited in what intelligence and data they can share with each other because of legal impediments around sharing SMRs between reporting entities to minimise the risk of (accidental) tipping off.
-There is also not a lot of feedback loops built into the system, including that between reporting entities and an FIU, and that between an FIU and law-enforcement agencies.
-The important limitations in the current state can thus be mostly traced to the following two issues:
-  • Complex financial crimes typically involve transactions that involve multiple financial institutions in multiple geographies, going through multiple payment channels, some of which are unregulated.
-  • FIUs and reporting entities are all working with data silos and no one has a complete view of activities across the financial system.
-To appreciate the extent of the problem, one only needs to look at the difficulty of trying to answer a simple question like finding all entities that exhibit a certain cross-border transaction pattern like that shown in Figure 3.
-Figure 3: A simple hard problem
-The question, and many more complicated others we would like to ask, simply cannot be answered today because the data needed to answer them sit in different organisations in different countries, governed by different legislation.
-2.2 Privacy Technologies As A Way Forward
-So what can we do?
-Figure 4 shows a possible way forward using distributed confidential computing technologies, assuming FIUs and partner agencies are willing to work together but without needing to make wholesale changes to AML/CTF legislations everywhere.
+The overall system is not perfect by any means.
+Nevertheless the regime is reasonably effective in deterring, detecting, and disrupting financial crimes, as demonstrated by the many successful prosecutions listed in Section 1.
+The main limitation of the current state is in the links that are there in the physical financial system but remain missing in the financial intelligence ecosystem.
+To begin with, a lot of financial transaction data, including domestic interbank transfers and crypto currency transactions, are not reported by default to an FIU and that is a major blind spot for analysts.
+Compounding the issue, the many thousands of reporting entities in a mature economy like Australia are heavily handicapped in their ability to share intelligence and data with each other because of legal impediments and risks around accidental tip-offs.
+One can also argue that there is simply not enough feedback loops built into the system, including that between reporting entities and an FIU, and that between an FIU and its partner law-enforcement agencies and foreign FIUs.
+In summary, the key limitations in the current state of financial intelligence can be traced to the following issues:
+  • 1.
+    Sophisticated and complex financial crimes typically involve transactions that span multiple financial institutions in multiple geographies, going through multiple payment channels, some of which remain unregulated.
+  • 2.
+    Criminality cannot be established through analysis of financial transaction data alone; actual links to predicate crimes are necessary.
+  • 3.
+    FIUs, reporting entities, and law-enforcement agencies are all mostly working within data silos and it is hard for any one organisation to get a complete view of financial and criminal activities across the entire system.
+2.2 A Way Forward using Privacy Technologies
+To overcome the key limitations of the current state of financial intelligence, Figure 3 outlines a potential solution that leverages distributed confidential computing technologies.
+The assumption is that FIUs, reporting entities and law-enforcement agencies are willing to collaborate in good faith without requiring sweeping changes to AML/CTF legislation globally.
 Figure 3: Privacy-enhanced version of Figure 2. LEAs = Law Enforcement Agencies; PAs = Partner Agencies; EOI = Entities of Interest
-The obvious obstacle to different FIUs and partner agencies sharing large volumes of data with each other is privacy concerns.
-In the proposed scheme, instead of entity resolution, we have all the participating agencies contributing data that are first matched using privacy-preserving record linkage (PPRL) algorithms [CRS20].
-In a PPRL algorithm, the records in different databases that represent or belong to the same entity are identified in a privacy preserving manner without revealing any plaintext values to the participating parties.
-The matched data are then encrypted using a homomorphic encryption scheme [Hal17] (as described in Section 3.2), which allow mathematical operations to be performed on encrypted data, to arrive at a table similar to what we showed earlier, with two distinctions:
-  • The values in the table are all encrypted and the columns of the table are stored in a distributed way across different databases in different agencies.
-  • The point of using homomorphic encryption schemes is that it is still possible to run a range of risk-modelling algorithms over the encrypted data, and once we get to a small set of entities of interest, decrypting those entities’ data can be done without raising any privacy concerns for the general public.
-Having the data distributed in different data sources does complicate the running of risk-modelling algorithms somewhat but the added security and scalability make it a worthwhile trade-off to make.
-There are basically three fundamental computational problems to solve, in increasing order of difficulty:
+In the proposed approach, rather than traditional entity resolution, data from all participating organisations are first matched using privacy-preserving record linkage (PPRL) algorithms [18].
+These algorithms identify records across different databases that correspond to the same entity in a manner that safeguards privacy, ensuring no plaintext values are disclosed to any party involved.
+Once the records are matched, the data are encrypted using either homomorphic encryption (Section 3.2) or secure multiparty computation schemes (Section 3.1).
+These encryption methods allow mathematical operations to be performed directly on encrypted data, resulting in a table similar to those discussed earlier but with two key differences:
+  • 1.
+    All values in the table are encrypted, and the columns are distributed across multiple databases managed by different agencies.
+  • 2.
+    AI and machine learning algorithms for detecting risks and suspicious entities can still be run on the encrypted data using homomorphic encryption or secure multiparty computation.
+When a small set of entities of interest is identified, their data can be decrypted without compromising the privacy of the broader population.
+Distributing encrypted data across multiple organisations and databases does introduce some complexity to AML/CTF risk-modelling workflows, but the approach offers a pathway for collaboration while addressing privacy concerns and ensuring compliance with existing legal frameworks.
+To help us navigate the opportunities offered by privacy technologies for financial intelligence, it is essential to understand that there are three fundamental computational problems to address in AML/CTF risk modelling, in increasing order of technical complexity:
   1. 1.
-    How can two FIUs, each of which has a list of suspicious entities, share only the common suspicious entities with each other, without sharing their entire list with each other?
-More generally, how can such sharing be done between multiple FIUs?
+    Private Intersection Identification: How can two FIUs, each maintaining its own list of suspicious entities, securely exchange overlapping data points without exposing their full datasets with each other?
+How might this process scale across multiple FIUs?
   2. 2.
-    Given a criminal typology that can be described as a graph pattern, where nodes denote entities and edges denote relationships, how can we detect all entities that satisfy the criminal typology when the underlying graph span multiple financial institutions?
+    Cross-Institutional Pattern Detection: Given a criminal typology defined as a graph pattern (where nodes represent entities and edges denote relationships), how can we identify all entities matching this pattern when the underlying graph pattern spans multiple financial institutions?
   3. 3.
-    How can we perform model learning, either supervised or unsupervised, over a set of entities whose financial data and behaviour are spread across multiple financial institutions?
-In the rest of this paper, we will look at how aspects of that future state can be realised through different privacy technologies.
+    Federated Model Learning: How can supervised or unsupervised machine learning models be trained on behavioural and transactional data distributed across multiple financial institutions, where no single entity holds complete datasets?
+These three computational problems are distilled from the authors’ experience working both directly and indirectly with Australia’s financial intelligence and banking communities over an extended period of time, and backed up by literature review.
+In the rest of this paper, we will first survey some key privacy technologies in § 3 before providing in § 4 illustrative examples and detailed pointers to how such privacy technologies can be used to address the three computational problems listed above.
+Solutions to these three computational problems can be, and are often, used as building blocks to develop more complete systems to address challenging real-world financial intelligence problems, some of which we survey in § 5.
 3 Privacy Technologies
-The purpose of this section is to briefly introduce the most common privacy technologies that are applicable to financial intelligence.
-These techniques are based on different mathematical foundations to serve different data privacy and security purposes.
-Secure multiparty computation, through the use of secret sharing schemes, permits multiple parties to jointly compute a function without divulging each party’s secret information to the others.
-Homomorphic encryption tackles this issue from a different perspective, by encrypting sensitive data in a way that allows arithmetic operations to be performed directly on the ciphertexts.
-Some privacy issues could arise through reverse-engineering the algorithm or query outcomes [DN10, KM11].
-To address this, the differential privacy framework can be used to provide individuals with guaranteed plausible deniability by adding suitably calibrated noise to query outcomes.
-We will also describe federated learning, which is needed for multiple parties to jointly learn a machine learning model from distributed data.
-We will start the section with hashing, a foundational technique for many privacy technologies.
+In the above, we highlighted a central challenge in multiparty financial intelligence, how can multiple parties work together effectively when none of them can be assumed to be fully trusted?
+We also noted that privacy technologies can serve as a technological means to enable such collaboration, allowing multiple parties to jointly combat financial crime under specific cooperation models.
+We refer to this form of multiparty financial intelligence conducted under explicit data privacy and security constraints as privacy-preserving financial intelligence.
+Privacy-preserving financial intelligence should be understood as a system level concept rather than a single method.
+Different privacy technologies play different roles within this system, but they share a common objective of enabling analysis and modelling, while maintaining data security and privacy guarantees.
+This section begins with secure multiparty computation (MPC), as it is one of the most widely adopted mechanisms for protecting data in collaborative settings.
+MPC enables multiple parties to jointly perform computations without revealing their private inputs.
+Building on this foundation, we introduce homomorphic encryption (HE), which supports MPC by enabling computation over encrypted data and is therefore often used as a cryptographic primitive in MPC designs.
+We then move beyond MPC centric settings and discuss several problem scenarios where MPC is not the primary solution.
+For these cases, we introduce privacy technologies that are widely used in practice, namely hashing, federated learning (FL), and differential privacy (DP), each addressing privacy risks from a different perspective.
+3.1 Secure Multiparty Computation
+The goal of secure multiparty computation (MPC) is to enable multiple parties to jointly evaluate a pre-agreed function without revealing their individual private inputs.
+At its core, MPC removes the need to trust the computation process itself.
+Under an MPC protocol, each participant learns only the final output (or an authorised share of it) and cannot infer additional information about other parties’ inputs beyond what is inherently revealed by that output.
+In doing so, MPC shifts cross-institutional collaboration away from data centralisation or reliance on a trusted third party, and instead enforces privacy and correctness directly at the protocol level through cryptographic mechanisms.
+More generally, MPC can be viewed as a general-purpose framework for privacy-preserving computation.
+Whenever a collaborative task can be formalised as a function, MPC provides a mechanism for jointly evaluating that function without requiring participants to share their raw data.
+Because MPC is a strict security framework, its practical deployment depends on clearly defined security guarantees.
+In particular, privacy ensures that an adversary learns nothing about an honest party’s input beyond what follows from the output itself.
+Correctness guarantees that the protocol produces the correct result, even in the presence of adversarial behaviour.
+Stronger notions may also be required in certain settings: guaranteed output (or robustness) ensures that adversaries cannot prevent honest parties from obtaining the result by aborting the protocol, while fairness requires that an adversary receives the output if and only if the honest parties do.
+In addition to these properties, MPC protocols are analysed under explicit adversarial models.
+In the semi-honest (honest-but-curious) model, adversaries follow the protocol specification but attempt to extract additional information from the execution.
+In the malicious model, adversaries may deviate arbitrarily from the protocol, for example by sending malformed messages, aborting prematurely, or manipulating inputs and outputs.
+The assumed adversarial model directly influences both the strength of the security guarantees and the complexity of the resulting protocol.
+In the context of financial intelligence, these capabilities can be illustrated through a common investigative scenario (discussed further in Section 4.1).
+Different financial institutions (FIs) maintain their own customer lists, while a financial intelligence unit (FIU) may need to determine which entities on its watchlist have relationships with which banks.
+Here, the FIU and the banks act as participating parties, their private inputs are their respective customer lists, and the agreed computation is the intersection between the FIU’s list and those of the banks.
+MPC enables this joint analysis to be conducted without exposing any party’s complete customer list.
+3.2 Homomorphic Encryption
+Within the MPC framework, privacy protection is not achieved by simply hiding the output, but by ensuring that computation itself can be performed without exposing sensitive inputs.
+For this reason, MPC protocols typically rely on multiple cryptographic mechanisms rather than a single technique.
+Homomorphic encryption (HE) is one such foundational capability.
+The defining feature of HE is that it allows certain computations to be performed directly on encrypted data, producing results that, once decrypted, are equivalent to those obtained from computations on plaintext.
+In MPC settings, this property enables parts of a computation to be executed without revealing intermediate values, thereby reducing direct data exposure between participants.
+$\text{Dec}(\text{secret key},\text{Enc}(\text{public key},m_{1})\circ\text{Enc}(\text{public key},m_{2}))=m_{1}\diamond m_{2},$ (1)
+From this perspective, HE is not a replacement for MPC, but rather one of its supporting building blocks.
+In practice, HE is often combined with other mechanisms, such as secret sharing or oblivious transfer, to securely encapsulate specific computation steps while balancing privacy guarantees, communication overhead, and system complexity.
+At the same time, HE can also be used independently of a full MPC framework.
+In scenarios where the computation structure is relatively simple, for example, when one party holds the data and another performs the computation, HE allows encrypted data to be processed externally without revealing the underlying inputs.
+This enables analysis or evaluation without requiring raw data sharing.
+HE schemes are commonly classified according to the types of operations they support and the number of operations that can be performed before decryption becomes unreliable.
+The most general form is Fully Homomorphic Encryption (FHE), which supports an unbounded number of operations on encrypted data.
+The practical feasibility of FHE was established following Gentry’s breakthrough construction [27], which sparked significant research interest, particularly in lattice-based cryptography, a field now widely regarded as central to quantum-resistant encryption [2].
+It is also important to note that HE schemes are inherently probabilistic, which means encrypting the same plaintext multiple times produces different ciphertexts.
+This property ensures semantic security, which prevents adversaries from inferring information about the plaintext from ciphertext patterns.
+In financial intelligence applications, such standalone use of HE may arise in outsourced computation of sensitive transaction features, risk metrics, or aggregate statistics.
+Financial institutions can retain full control over their data while leveraging external compute resources or cross-organisational collaboration, without disclosing plaintext information.
+Despite its strong privacy guarantees, the applicability of HE remains constrained by performance considerations, supported computation types, and overall system complexity.
+As such, HE is best understood either as a specialised independent capability within the privacy-preserving toolkit, or as a key component within more complex MPC-based solutions.
 3.3 Hashing Techniques
-Hashing has a long history in Computer Science.
-Hashing was first introduced in 1953 to speed up the searching for information in storage systems [Kno75].
-The general idea of hashing, or more specifically a hash function, is to map values that are of arbitrary sizes, such as textual strings that can have different lengths, to fixed sized values such as integer numbers that are in a specific range.
-A hash function therefore converts input values into hash values or hash codes.
-These hash values are then used as indexes or pointers into a hash table of fixed size.
-According to [CZ17], a good hash function needs to: (1) map values to hash values as uniformly as possible across the space of possible hash values, (2) the same input value will always result in the same hash value, (3) the mapping process should be fast, and (4) the size of the hash table for a given set of values can be specified by setting the range of integer hash values.
-Hashing has been and continues to be a valuable security mechanism for making data unreadable to the human eye, preventing its interception by malicious individuals, and providing a way to validate its integrity.
-Next we explain few variations of hashing that are useful in the financial intelligence domain.
-Keyed Cryptographic Hashing
-A keyed hash function (also known as a hash message authentication code, or HMAC) is an algorithm that uses a cryptographic key and a cryptographic hash function to produce a message authentication code that is keyed and hashed [BCK96].
-In general HMAC is used to verify the integrity and authenticity of data by computing a fixed-size hash value based on a secret key and input data.
-Unlike traditional cryptographic hash functions that generate a hash value solely based on the input data, keyed hashing incorporates a secret key or salt value into the hashing process, making it resistant to tampering and unauthorised modification.
-The key serves as a shared secret between the sender and receiver, ensuring that only parties possessing the key can compute or verify the hash value.
-This provides a strong level of security against malicious tampering or forgery attempts, as any alteration to the input data or hash value would require knowledge of the secret key to generate a valid HMAC.
-Keyed cryptographic hashing is widely used in various security applications, including message authentication, data integrity verification, and secure communication protocols [KSN+23].
-Bloom Filters
-A Bloom filter [Blo70] is a space-efficient probabilistic data structure used to test whether an element is a member of a set.
-It is implemented as a bit array $BF[i]$, where $i\in[1,n]$, consisting of $n$ bits.
-The Bloom filter uses $k$ independent hash functions $H_{1}(x),\cdots,H_{k}(x)$, each of which maps an element $x$ in a set $S$ of $m$ elements to one of the $n$ bits within the bit array $BF$.
-To determine if an element $x$ is present in the Bloom filter, one needs to check if all $k$ hash functions map $x$ to bits that are set to 1 in the Bloom filter.
-Because of hash collisions, false positives can occur in a Bloom filter, but false negatives are impossible.
-Figure 5: The Dice similarity calculation between the names ‘peter’ and ‘pete’, converted into bigrams ($q=2$) and then encoded into two Bloom filters $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ of length $l=12$ bits using $k=2$ hash functions.
-The bit “$1$” shown in italics at position $6$ in $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ is a hash collision, because both ‘pe’ and ‘te’ are hashed to this position.
-In the privacy-preserving record linkage domain, Bloom filters are commonly used for matching string values or computing similarities between string pairs in a privacy-preserving manner [SBR09, CRS20].
-When matching string values, each string is first converted into a set of q-grams, where each q-gram is typically of length 2 or 3.
-These q-grams are then encoded into a Bloom filter using $k$ hash functions.
-To compute similarities between Bloom filters, a set-based similarity function such as Jaccard or Dice can be used.
-Recent literature has also shown that Bloom filters can be used to match both numerical and textual values in databases [WVKR23].
-Figure 5 illustrates an example of using a Bloom filter for data matching.
-The Privacy-Preserving Record Linkage step in Figure 4 can be solved using variations of Bloom Filters.
-The hash functions in such Bloom Filters are usually keyed cryptographic hashing functions to protect against dictionary attacks.
-3.2 Homomorphic Encryption
-Homomorphic encryption (HE) is a class of encryption schemes that allows certain arithmetic operations to be performed directly on the ciphertext [AAUC18].
-HE schemes are usually public-key cryptosystems.
-Let $m_{1}$ and $m_{2}$ be two plaintexts, and let $pk$ and $sk$ be the public key and secret key for the encryption Enc() and decryption Dec() functions, respectively.
-The correctness of a HE scheme is expressed by the following property.
-For an operation $\diamond$ in the plaintext space, there is a corresponding operation $\bullet$ in the ciphertext space such that
-$\text{Dec}(sk,\text{Enc}(pk,m_{1})\bullet\text{Enc}(pk,m_{2}))=m_{1}\diamond m_{2},$ (1)
-where the different symbols indicate that the arithmetic operations, either addition or multiplication, in the plaintext and ciphertext spaces may differ.
-The term “homomorphic” reflects that the decryption function acts like a homomorphism, preserving the structure of the mathematical object.
-Depending on the types of arithmetic operations permitted and the number of operations it can handle before decryption fails, an HE scheme is classified into one of the following types, where the algorithmic complexity of the operations we want to perform on data is measured in terms of the depth of circuits and the arithmetic gates supported.
-  • Partially HE (PHE) - Schemes that can evaluate circuits containing only one type of arithmetic gate, that is, either addition or multiplication, for unbounded circuit depth.
-  • Somewhat HE (SHE) - Schemes that can evaluate a subset of circuits containing both addition and multiplication gates, where the complexity grows as the circuit depth increases.
-  • Fully HE (FHE) - Schemes that can evaluate all circuits containing both addition and multiplication gates for unbounded circuit depth.
-The privacy-preserving record linkage and model scoring steps in Figure 4 can be done using PHE, but full-scale risk modelling using both supervised and unsupervised machine learning algorithms would usually require SHE or even FHE.
-FHE was made possible only in recent years due to the breakthrough described in [Gen09].
-Gentry’s work is technically based on lattice-based cryptography, an active research area also seen as a strong candidate for quantum-resistant encryption [A+22].
-In addition to the standard HE components, Gentry’s scheme introduces two key concepts: squashing and bootstrapping.
-Simply put, squashing reduces the complexity of decryption by providing a hint to the evaluation process (which introduces additional security assumptions), while bootstrapping allows the scheme to evaluate its own decryption along with an arithmetic operation.
-When combined, these techniques enable the scheme to perform unbounded arithmetic operations with controlled decryption errors.
-It is worth noting that HE schemes are inherently probabilistic, meaning that the same plaintext can produce a different ciphertext each time it is encrypted.
-This is to ensure a security notion called semantic security.
-For a more thorough introduction to HE and its mathematical and computer science foundations, the reader is referred to [Hal17, Pei16, LNP22].
-To conclude this section, we provide an example of a partially homomorphic encryption (PHE) scheme: the multiplicative HE scheme known as ElGamal Encryption [ElG85].
-This scheme is based on the Diffie-Hellman key exchange protocol [DH76] to securely exchange a masked secret key between Alice and Bob for encryption and decryption.
-Public key ElGamal Encryption. Bob is the sender, Alice is the receiver. Alice chooses a cyclic group $G$ with order $q$ and a generator $g$.
-Private key: Alice’s secret key is a random integer $x\leftarrow\{1,\dots,q-1\}$.
-Public key: Alice computes $h=g^{x}\in G$ and publishes the public key $(G,q,g,h)$.
-Encryption: To encrypt a message $m\in G$, Bob randomly selects $y\leftarrow\{1,\dots,p-1\}$, computes the shared secret $s=h^{y}\in G$, and sends the ciphertext $(c_{1},c_{2})$, where $c_{1}=g^{y}$ and $c_{2}=m\cdot s$.
-Decryption: To decrypt, Alice computes the shared secret $s=c_{1}^{x}$ and its inverse $s^{-1}\in G$, then recovers the message $m=c_{2}\cdot s^{-1}$.
-Since each term in the ElGamal ciphertext can be multiplied separately to produce a multiplication ciphertext, and Alice can recover the correct multiplied message using the above decryption process, it implies that ElGamal is a multiplicative HE scheme.
-Specifically, if $C=(c_{1},c_{2})\leftarrow\text{Enc}(m)$ and $\tilde{C}=(\tilde{c}_{1},\tilde{c}_{2})\leftarrow\text{Enc}(\tilde{m})$, then decryption of the product $C\cdot\tilde{C}$ yields $m\cdot\tilde{m}$, i.e., Dec$(C\cdot\tilde{C})=m\cdot\tilde{m}$.
-We will see some applications of the additive homomorphic form of ElGamal encryption in Section 4.2.
-3.3 Secure Multiparty Computation
-Secure multiparty computation (MPC) is a concept that drives the development of protocols designed to securely and accurately perform computations across multiple parties.
-Research in MPC originated with Yao’s so-called Millionaires’ problem [Yao82], in which two parties seek to know who is richer but without having to disclose their net wealth to each other or a trusted third party.
-There are many solutions to the Millionaires’ problem involving different cryptographic primitives.
-Here we briefly describe an accessible solution from [LT05].
-The algorithm first converts two numbers in binary representation $x=x_{n}x_{n-1}\ldots x_{1}$ and $y=y_{n}y_{n-1}\ldots y_{1}$ we want to privately compare to the following two sets
-$S_{x}=\{x_{n}x_{n-1}\ldots x_{i+1}1\,|\,x_{i}=0,1\leq i\leq n\}$
-$S_{y}=\{y_{n}y_{n-1}\ldots y_{i}\,|\,y_{i}=1,1\leq i\leq n\}$
-and then securely compute the cardinality $|S_{x}\cap S_{y}|$.
-The correctness of the procedure comes from the fact that $x>y$ if and only if $S_{x}\cap S_{y}$ is non-empty.
-There are various private set intersection cardinality (PSI-CA) protocols for computing $|S_{x}\cap S_{y}|$, utilising different techniques such as homomorphic encryption, hashing, Bloom filters, and oblivious transfer.
-Here, we describe a PSI-CA protocol that employs hashing and commutative encryption from [AES03].
-Informally, an encryption function $f$ is commutative if the order of encryption does not matter for arbitrary encryption keys $e$ and $e^{\prime}$, i.e., $f_{e}\circ f_{e^{\prime}}=f_{e^{\prime}}\circ f_{e}$ where $\circ$ is function composition.
-An example given in [AES03] is the function
-$f_{e}(x)=x^{e}\mod p,$
-where the powers commute, yielding
-$(x^{e}\bmod p)^{e^{\prime}}\bmod p=x^{ee^{\prime}}\bmod p=(x^{e^{\prime}}\bmod p)^{e}\bmod p.$
-The security of this commutative encryption is based on the Decisional Diffie-Hellman hypothesis [Bon98].
-Armed with a commutative encryption function, $|S_{x}\cap S_{y}|$ can be computed via double encryption as follows:
-$|S_{x}\cap S_{y}|=|f_{k_{y}}(f_{k_{x}}(S_{x}))\cap f_{k_{x}}(f_{k_{y}}(S_{y}))|,$
-where $k_{x}$ and $k_{y}$ are the encryption keys for the two parties holding the value $x$ and $y$, respectively.
-The above secure number comparison protocol provides a glimpse into the rich literature on MPC [Gol98, CD15, EKR+18].
-Interestingly, what makes a distributed computation task an MPC protocol cannot be straightforwardly defined in a mathematical way.
-The well-accepted way of defining MPC’s security is by simulating an ideal world and a real world, known as the ideal/real simulation paradigm [Lin20].
-In the simulation, the two worlds are identical except that there is a trusted third party in the ideal world but not in the real world.
-To examine the security of a protocol under adversarial attacks, one would compare the outcomes of the protocol in the two worlds to determine if the adversary in the real world is able to gather any additional non-disclosed information, should the same attacks also be successful in the ideal world.
-With the ideal/real simulation paradigm in place, the following are important properties of an MPC protocol.
-Privacy guarantees that an adversary learns nothing about an honest party’s input beyond what is prescribed.
-Correctness ensures that the protocol always produces the correct result despite being attacked by adversaries.
-Guaranteed output stops the adversary from halting the protocol hence stopping others from receiving an output.
-Lastly fairness states that an adversary receives the protocol output if and only if an honest party receives the protocol output.
-There are two primary adversarial models: semi-honest and malicious adversaries.
-In the semi-honest model, an adversary correctly follows the protocol specification, but seeks opportunities to learn the protected privacy from information obtained during the computation process.
-In contrast, the malicious adversarial behaviour can arbitrarily deviate from the protocol, including aborting the protocol or changing its inputs and outputs.
+In cross-institutional data collaboration, a fundamental but often overlooked problem is how to preserve matchability between entities or relationships while avoiding direct exposure of sensitive identifiers.
+At this level, hashing provides a lightweight yet practical solution.
+Hashing works by applying a one-way mapping to original identifiers such as customer IDs, transforming them into fixed length representations.
+For example,
+$\text{Hash: Hello}\rightarrow\text{2CF24DBA5FB0A30E26E83B2AC5B9E}$
+While the original identifiers cannot be recovered, consistent hashing allows entities or relationships across different datasets to be matched.
+In other words, hashing does not hide whether two records refer to the same entity, it hides who that entity is.
+From a functional standpoint, hashing is not a joint computation mechanism and does not attempt to provide comprehensive protocol level privacy guarantees.
+Its primary value lies in reducing direct exposure of sensitive identifiers during data exchange and comparison, thereby offering a baseline level of privacy protection.
+This makes hashing widely used in practice for identity alignment, relationship linking, and preliminary data filtering.
+In financial intelligence contexts, hashing is commonly applied as a preprocessing step for customer or account identifiers.
+This allows financial institutions and regulators or investigators to assess potential overlaps or business relationships without sharing plaintext identity information.
+While such approaches do not eliminate all privacy risks, they can significantly reduce unnecessary data exposure when computation goals are well defined and trust boundaries are controlled.
+As suggested by Chi et al.
+[17], properties such as uniform distribution of hash outputs, consistency of mapping, computational efficiency, and appropriate handling of input and output domains are important when selecting a hash function for practical applications.
+More broadly, the effectiveness of hashing depends on several underlying assumptions, including the size of the input space, the consistency of hashing rules across participants, and resilience against auxiliary information attacks.
+In higher-risk or more complex multiparty collaboration settings, hashing alone may be insufficient; it is therefore typically combined with stronger privacy-preserving mechanisms rather than deployed as a standalone solution.
+3.4 Federated Learning
+In some collaborative settings, privacy risks arise not from a single joint computation, but from long term, iterative model training processes.
+In such cases, participants seek to collaboratively build and update models over time while keeping their data within local domains, rather than jointly evaluating a fixed function.
+Federated learning (FL) is designed for precisely this type of collaboration.
+The core idea of federated learning is that data remains local to each participant, model training is performed independently on local datasets, and only model updates are shared with a coordinating entity for aggregation into a global model as shown in Figure 4.
+From a collaboration perspective, FL does not require sharing raw data or encrypting an entire joint computation.
+Instead, it follows a local computation and global aggregation paradigm.
+Figure 4: An example federated learning (FL) architecture with $n$ data owners.
+In the setting, data owners only exchange the model gradient ($w$) with the aggregation server.
+In each training iteration, each data owner $i$ trains its own model with their local data and send the corresponding local model updates $w_{i}$ to the aggregation server.
+After aggregating all the received local model updates, the aggregations server returns the new global model updates ($W$) to each data owner.
+This leads to a fundamental structural difference between FL and MPC.
+While MPC focuses on securely evaluating a predefined function, FL is concerned with the continuous evolution of a model over time.
+In large scale, long running, and frequently updated training tasks, attempting to replicate FL using MPC would be impractical due to excessive system complexity, communication overhead, and reduced training efficiency.
+It should be noted, however, that although FL avoids direct data sharing, privacy risks do not disappear entirely.
+Model parameters or gradient updates may, under certain conditions, leak statistical properties or even individual level information about local datasets [3, 46, 66].
+Consequently, FL is often combined with additional privacy technologies, such as homomorphic encryption or differential privacy, to mitigate information leakage during the training process.
+Within financial intelligence, FL is particularly well suited to scenarios where multiple institutions wish to jointly train models for risk detection, anomaly identification, or behavioural analysis without centralising sensitive data.
+Participants retain control over their data while benefiting from improved model performance through collaboration.
+In such settings, the key challenge is not how to securely compute a single result, but how to balance model utility and privacy risk over sustained collaboration, a challenge for which FL offers distinct advantages over MPC.
 3.5 Differential Privacy
-Differential privacy (DP) is a technique that aims to enable the analysis of sensitive data while protecting the privacy of individuals whose data is included in the analysis [DR14].
-In general, DP is achieved by adding noise in a controlled manner during the analysis process, ensuring that statistical queries do not reveal information about specific individuals.
-This allows organisations to derive valuable insights from their data while minimising the risk of unauthorised disclosure or misuse of sensitive personal data.
-Definition 3.4.1.
-A randomised algorithm $\mathcal{A}$ is $\epsilon$-differentially private if for all neighbouring databases $\mathbf{D}$ and $\mathbf{D^{\prime}}$ that differ in at most one record, and for all sets $\mathcal{O}$ outputs, we have
+Across the privacy technologies discussed so far, the primary focus has been on protecting input data and ensuring trustworthy computation.
+However, in many real world applications, the results themselves can still leak sensitive information, even when inputs and computation processes are fully protected.
+Differential privacy (DP) addresses this class of risks directly.
+The core principle of DP is to introduce carefully calibrated randomness into statistical outputs or model results, ensuring that the presence or absence of any single individual’s data does not have a distinguishable impact on the output.
+Definition 1
+A randomised algorithm $\mathcal{A}$ is $\epsilon$-differentially private if for all databases $\mathbf{D}$ and $\mathbf{D^{\prime}}$ that differ in at most one record, and for all sets $\mathcal{O}$ outputs, we have
 $Pr[\mathcal{A}(\mathbf{D})\in\mathcal{O}]\leq\exp(\epsilon)\cdot Pr[\mathcal{A}(\mathbf{D^{\prime}})\in\mathcal{O}],$
 where $Pr[\cdot]$ denotes the probability of an event.
-The parameter $\epsilon$ is known as the privacy budget in DP.
-A smaller value of $\epsilon$ indicates a stronger privacy guarantee, as it reflects a stricter bound on how much the query results can change in response to the inclusion or exclusion of a single record.
-The Bayesian semantics of differential privacy can be found in [KS14, KM14, YZN24]; in short, DP guarantees that the odds ratio of an attacker’s prior and posterior beliefs (before and after seeing the query output) in whether a specific record is part of the input database is bounded, and that confers every individual with plausible deniability on whether the person’s data is in the dataset.
-An example scenario that illustrates Definition 3.4.1 is differencing attacks.
-Consider a medical database containing records of individuals who participated in a disease screening.
-An analyst with access only to counting queries can query the database under different conditions to select two distinct subsets of the data.
-If the difference between these subsets is designed to isolate a specific individual, the analyst could potentially infer whether or not this individual has the disease by examining the difference between the two query results.
-Differential privacy prevents this by injecting suitably calibrated noise into the results of these queries, thus preventing the analyst from reliably determining the presence or absence of any particular individual based on the differences in query results.
-A fundamental DP technique is the Laplace Mechanism [DMNS06, DR14].
-Given a real-valued query function $f:\mathbb{N}^{|\chi|}\rightarrow\mathbb{R}^{k}$, the Laplace Mechanism directly adds Laplace noise to the query outputs as
-$M_{L}(\mathbf{D},f,\epsilon)=f(\mathbf{D})+(e_{1},\dots,e_{k}),\text{ where }e_{i}\leftarrow\text{Lap}\left(\frac{\Delta f}{\epsilon}\right),$
-where Lap() is the Laplace distribution, and
-$\Delta f=\max_{\mathbf{D}\sim\mathbf{D}^{\prime}\in\mathbb{N}^{|\chi|}}{||f(\mathbf{D})-f(\mathbf{D}^{\prime})||_{1}}$
-is a measure of the sensitivity of the query on two neighbouring databases $\mathbf{D}$ and $\mathbf{D}^{\prime}$.
-The independent and identically distributed noise is randomly sampled from the Laplace distribution with mean 0 and scale $\Delta f/\epsilon$.
-The noise’s scale is proportional to the query sensitivity (when $\epsilon$ is fixed) and inversely proportional to the noise budget $\epsilon$ (when $\Delta f$ is fixed).
-This makes good sense, since a less sensitive query needs less noise to hide its output variation.
-A smaller $\epsilon$ indicates stronger privacy guarantee hence more noise needs to be added to the outputs.
-In the financial domain, where large volumes of sensitive data are collected and analysed for various purposes such as risk assessment, fraud detection, customer segmentation, and personalised marketing, DP serves as a valuable tool for balancing the need for data-driven insights with privacy protection.
-By incorporating DP into data analytics workflows, financial institutions can adhere to stringent regulatory requirements, such as GDPR and Basel III [^5] See https://gdpr-info.eu and https://www.bis.org/bcbs/basel3.htm, while still leveraging the richness of their datasets for business intelligence and decision-making [Nis21].
-Essentially all financial data are graph data and a survey of DP algorithms for graph data can be found in [LPR+23].
-3.5 Federated Learning
-Federated learning (FL) [MMR+17] is a new learning paradigm that aims to build a joint machine learning model based on data distributed across multiple sites or owned by different participants.
-During model training, information such as gradients is exchanged between participants, but the raw data are not shared.
-The intention is that the exchanged information does not reveal any private data belonging to any party.
-Figure 6: A federated learning architecture.
-In order to guarantee the privacy of the data, federated learning only permits all the data owners to exchange the model gradient ($w$) with the aggregation server.
-During this process, each data owner trains its own model with the local data, then uploads the local model to the aggregation server.
-After aggregating all the received models, the server returns the new global model updates ($W$) to each data owner.
-Figure 6 illustrates a typical FL process.
-Initially, each data owner receives a generic global model from the aggregation server.
-After receiving the initial model, each data owner conducts local training on this model with their own data and then uploads the related gradient information (local model updates) to the aggregation server.
-The aggregation server averages the updates received from data owners into the global model, then updates the global model to replace each user’s local model.
-This process repeats until the global model achieves the desired performance or the training reaches the maximum iteration number.
-FL can be categorised into three types based on the distribution of data: horizontal FL, vertical FL, and federated transfer learning [LSTS20].
-Horizontal FL occurs when the user features of the datasets overlap significantly, but the users themselves overlap minimally.
-Vertical FL occurs when the user features of the datasets overlap minimally, but the users themselves overlap significantly.
-In cases where both users and user features of the datasets overlap rarely, transfer learning can be used to address the lack of training data [LKX+20].
-Privacy protection is a major concern in FL [ARPS20, NDP+21].
-Although FL protects private data on each device by exchanging only model gradients with the aggregation server, the model communication during the entire training process can still potentially leak sensitive information to third parties, for example, through reverse engineering of models [ZLD22].
-Two techniques can be employed to prevent private data leakage during FL model transmission:
-  • Secure aggregation techniques ensure that raw data or local updates remain private during the aggregation process.
-Methods such as MPC or HE (described in Sections 3.3 and 3.2) can be used to aggregate updates without revealing individual contributions.
-  • DP methods add noise to local updates to protect individual privacy.
-This noise ensures that the contribution of any single device is indistinguishable from the contributions of others.
-Aggregating noisy updates provides privacy guarantees while still allowing the central model to learn from the collective data.
-In federated learning, model scoring and model learning are two distinct phases of the process.
-Model scoring refers to the evaluation or assessment of the performance of a machine learning model using a set of evaluation metrics.
-Common evaluation metrics for model scoring include accuracy, precision, recall, F1-score, and area under the ROC curve (AUC).
-Both federated model learning and federated model scoring are important steps in the future-state financial intelligence proposed in Figure 4.
+Rather than hiding raw data or computation steps, DP explicitly constrains how much information can be inferred from released results.
+This positioning fundamentally distinguishes DP from MPC.
+Even in a fully trusted joint computation environment, repeated queries or downstream inference based on exact outputs can gradually expose individual level information.
+DP mitigates this risk by limiting the sensitivity of outputs to individual data points, something that MPC, HE, or FL cannot achieve on their own.
+In financial intelligence applications, DP is commonly used when publishing aggregate statistics, analytical reports, or model outputs.
+For example, when sharing risk distributions, trend analyses, or group level behavioural insights, DP mechanisms can preserve the overall analytical value while substantially reducing the risk of re-identifying individual customers or accounts.
+In these contexts, the primary privacy concern is not who performed the computation, but what can be inferred once the results are released or reused [47, 38].
+It is important to emphasise that differential privacy is not a zero cost solution.
+Introducing randomness to limit information leakage necessarily involves a trade-off between privacy guarantees and result accuracy.
+As such, DP is best understood as an output level privacy governance mechanism, complementing rather than replacing input level and computation level privacy technologies.
 4 Privacy-Preserving Information Sharing
-In this section, we describe several real-world financial intelligence scenarios wherein information-sharing between different parties is crucial to support effective detection of financial crimes.
+In this section, we revisit the three computational problems listed at the end of § 2 and further contextualise them using real-world financial intelligence scenarios wherein information-sharing between different parties is crucial to support effective detection of financial crimes.
 In each scenario, we discuss the privacy risks during multiparty collaborations and examine the feasibility of applying various privacy-preserving techniques to address these privacy risks.
@@
 Problem statement
-Suppose an FIU possesses a moderate-sized list of suspicious individuals (or, more generally, commercial entities) obtained, say, from its own investigations and law-enforcement partner agencies and the FIU wishes to acquire additional information about those suspicious individuals, including where they conduct their financial activities and whether they are known to other FIUs.
-This can be done if the FIU can safely distribute the Suspicious Person List (SPL) to other FIUs and financial institutions to verify the association of individuals on the list with specific financial entities before engaging in further targeted information requests or exchanges with those financial entities.
+Suppose an FIU possesses a moderate-sized list of suspicious individuals (or, more generally, commercial entities) obtained, say, from its own investigations and law-enforcement agencies and the FIU wishes to acquire additional information about those suspicious individuals, including where they conduct their financial activities and whether they are known to other FIUs.
+This can be done if the FIU can safely distribute the Suspicious Person List (SPL) to other FIUs and financial institutions to verify the association of individuals on the list with specific financial entities before engaging in further targeted information exchanges with those financial entities.
 Privacy risks
 If the SPL is shared with the financial institution in a manner that allows for the identification of individuals on the list, there arise the following privacy concerns, especially in the low-trust public-to-private information-sharing setting.
+  • 1.
     (Accidental Tip-off) The financial institution may inadvertently disclose the FIU’s suspicions to an existing customer, prompting the customer to take additional measures to conceal their activities.
+  • 2.
     (Debanking and Displacement) The financial institution may deliberately cease its banking relationship with a suspicious individual before conclusive evidence can be gathered to substantiate or refute the individual’s suspected wrongdoing.
@@
 Further, suspected individuals that are debanked will also move on to other parts of the financial system, including the shadow financial system, and the FIU will lose sight of them.
+  • 3.
     (Need to Know) The FIU can inadvertently disclose the identities of suspicious entities who are not customers of a financial institution, and that would be an unacceptable violation of those individual’s privacy and the need-to-know principle.
@@
 The problem described above of identifying common individuals between the SPL and each participating financial institution’s customer database can be formulated as a private set intersection (PSI) problem.
-In a two-party computation protocol, the goal of PSI is to identify the common elements that appear in both parties’ sets and nothing else.
-We now describe a simple semi-honest PSI protocol from [FNP04] that utilises polynomial evaluation and homomorphic encryption to ensure the protection of unmatched individuals in the SPL.
-For detailed security and efficiency analyses, readers are encouraged to refer to [FNP04].
+The goal of PSI is to identify the common elements that appear in the parties’ sets and nothing else.
+We now describe a simple semi-honest PSI protocol from [26] that utilises polynomial evaluation and homomorphic encryption to ensure the protection of unmatched individuals in the SPL.
+For detailed security and efficiency analyses, readers are encouraged to refer to [26].
 For simplicity, it is assumed that each individual can be uniquely identified by an identification (ID) number, which can be derived from a national ID (in some countries) or a combination of personal information.
@@
     which is uniquely defined according to the Fundamental Theorem of Algebra.
-The FIU then encrypts the coefficients of $f(x)$ using HE and send the encrypted coefficients and the public key to the other parties:
+The FIU then encrypts the coefficients of $f(x)$ using HE and sends the encrypted coefficients and the public key to the other parties:
     $\text{Enc}(f)=\{\text{Enc}(a_{0}),\dots,\text{Enc}(a_{n})\}.$
@@
   4. 4.
-    Once the matching outputs are shared, FIU decrypts the outputs and find those elements that also appear in its suspicious person list.
+    Once the matching outputs are shared, FIU decrypts the outputs and finds those elements that also appear in its suspicious person list.
 By executing this PSI protocol, the FIU and financial institutions collaborate in a confidential manner, enabling the FIU to identify individuals who appear on both the FIU’s SPL and the databases of participating financial institutions.
@@
 Industrial grade solutions
-The FIU.net operated jointly by FIUs in the European Union has offered privacy-preserving sharing of suspicious entities for over a decade now, using the Ma3tch algorithm [BM13].
-A recent legal review of FIU.net can be found in [Mou20].
-In Australia, the Australian Financial Crime Exchange [^6] http://afcx.com.au is a platform for financial institutions to share lists of entities of interest, from a financial crime and cyber crime perspective, with each other and, selectively, with the Australian Taxation Office.
-There are now many fast PSI algorithms that can operate at large scale, including [CLR17, RT21, CM20].
+The FIU.net operated jointly by FIUs in the European Union has offered privacy-preserving sharing of suspicious entities for over a decade now, using the Ma3tch algorithm [6].
+A recent legal review of FIU.net can be found in [45].
+In Australia, the Australian Financial Crime Exchange (http://afcx.com.au) is a platform for financial institutions to share lists of entities of interest, from a financial crime and cyber crime perspective, with each other and, selectively, with the Australian Taxation Office.
+There are now many fast PSI algorithms that can operate at large scale, including [16, 53, 15].
 A key issue with PSI is that an adversary can use repeated PSI computations to completely recover an organisation’s entire dataset.
-Proposals like [PLN23] seek to address this issue by introducing DP protection into PSI computations.
+Proposals like [50] seek to address this issue by introducing DP protection into PSI computations.
 4.2 Transaction Flow Tracing and Typology Matching
@@
 A key task in financial intelligence supported by transaction-flow tracing is the identification of accounts that match certain money-laundering or terrorism-financing (ML/TF) typologies.
-The NDIS fraud described in [BILC23] serves as an example.
+The NDIS fraud described in [10] serves as an example.
 Australia’s National Disability Insurance Scheme (NDIS) provides financial support to eligible participants to improve their quality of life.
-However, continuous fraudulent activities against the NDIS scheme have resulted in several billions of dollars in losses from its intended use [AUS20].
+However, continuous fraudulent activities against the NDIS scheme have resulted in several billions of dollars in losses from its intended use [4].
 These fraudulent transactions share common patterns, crossing multiple organisations and eventually being transferred overseas.
@@
 Proposed solution
-The FinTracer solution proposed in [BILC23] is basically a way to perform an efficient and privacy-preserving breadth-first search across a large distributed graph in homomorphically-encrypted space, allowing Financial Intelligence Units like AUSTRAC to work with law-enforcement agencies and financial institutions to start with some known suspicious entities and find their footprint across the financial system
+The FinTracer algorithm described in [10] is a computationally efficient way to conduct breadth-first search on large distributed graphs in homomorphically-encrypted space.
+The algorithm enables Financial Intelligence Units to start with some known suspicious entities and then work with law-enforcement agencies and financial institutions to find the entire footprints of those entities across the financial system
   1. 1.
-    without having to disclose the suspicious entities to financial institutions, thus avoiding the risk of accidentally tipping off the entities; and
+    without having to disclose the actual identities of the suspicious entities to financial institutions, thus avoiding the risk of accidentally tipping off the entities; and
   2. 2.
-    without having to violate the privacy of unrelated entities even though their data need to be processed.
-The key observation behind this algorithm is that large distributed graphs can be represented using sparse matrices and graph traversal can be expressed as matrix multiplication operations [Bur21].
-This basic graph search algorithm can serve as a building block to detect a large range of financial crime typologies (i.e.
-graph patterns), thus significantly advancing the state-of-the-art in intelligence agency’s ability to detect complex financial crimes in a responsible, privacy-preserving manner at a large scale.
-[…]
-Figure 7: A simplified transaction network to demonstrate FinTracer on the NDIS example.
-Accounts with cross organisation transactions are linked by directed edges.
+    without having to compromise the privacy of unrelated entities even though their data need to be processed.
+The key insight behind the FinTracer algorithm is that large distributed graphs can be represented as sparse matrices and the breadth-first search algorithm can be expressed as a sequence of matrix multiplication operations [12].
+The FinTracer algorithm can be used as a foundational routine to detect a wide range of financial crime typologies (i.e.
+graph patterns).
+Its adoption can thus meaningfully advance the state-of-the-art in financial intelligence agency’s ability to detect complex financial crimes that span multiple organisations in a privacy-preserving manner at scale.
+a1a2a3a4a1b1b2b3b4b5c1c2c3c4c5d1d2d3d4 AA $B$ $C$ $D$
+Figure 5: A simplified transaction network to demonstrate FinTracer.
+Accounts with cross-organisation transactions are linked by directed edges.
 There are four financial institutions A, B, C and D, each having a number of accounts.
-The account a1 in A directly received NDIS payment.
+The account $a_{1}$ in A directly received NDIS payment.
 To see how this works, suppose AUSTRAC is interested in identifying the accounts (from multiple financial institutions) that transferred NDIS funds overseas.
@@
 These tags are homomorphically encrypted to securely follow NDIS funds in the entire transaction network.
-The key components of the FinTracer process are the ElGamal encryption scheme [ElG85] over the additive Curve25519 for its performance efficiency [Ber06], and the aggregations of the tag values across financial institutions over multiple iterations of the breadth-first search algorithm.
+The key components of the FinTracer process are the ElGamal encryption scheme [24] over the additive Curve25519 for its performance efficiency [7], and the aggregations of the tag values across financial institutions over multiple iterations of the breadth-first search algorithm.
 Figure 5 is a simplified example of the NDIS fraud scenario.
 In this example, there are four financial institutions.
-The only account that directly received suspicious NDIS fund is a1 in institution A.
+The only account that directly received suspicious NDIS funds is a1 in institution A.
 All institutions are only aware of the source and destination accounts that are involved in the transactions they process.
-In its simplest form (Algorithms 1 and 2 [BILC23]), the FinTracer protocol does the following steps:
-  • Initialisation: Each financial institution sets the initial FinTracer tag of each of its customer account $x$ to either $\text{Enc}(1)$ or $\text{Enc}(0)$, depending on whether $x$ receive NDIS fund directly.
+In its simplest form (Algorithms 1 and 2 [10]), the FinTracer protocol does the following steps:
+  • 1.
+    Initialisation: Each financial institution sets the initial FinTracer tag of each of its customer account $x$ to either $\text{Enc}(1)$ or $\text{Enc}(0)$, depending on whether $x$ receive NDIS funds directly.
 In the example above,
+    • (a)
       $t_{\leq}(a1)=\text{Enc}(1)$ and $t_{\leq}(\{b1,b2,c1,c2,d1\})=\text{Enc}(0)$ (the set notation denotes that the tag is set for element in the set).
+    • (b)
       The other accounts do not have cross-institution transactions, hence are out of considerations by the protocol.
+  • 2.
     Iteration 1: In the first iteration, the initialised tags are passed to the neighbouring institutions following the transaction flow.
@@
 In the example, the shared partial mappings are
+    • (a)
       A $\rightarrow$ B: $P^{A}_{B}((a1,b1))=\text{Enc}(1),P^{A}_{B}((a1,b2))=\text{Enc}(1)$.
+    • (b)
       B $\rightarrow$ C: $P^{B}_{C}((b1,c1))=\text{Enc}(0),P^{B}_{C}((b2,c2))=\text{Enc}(0)$.
+    • (c)
       C $\rightarrow$ D: $P^{C}_{D}((c1,d1))=\text{Enc}(0),P^{C}_{D}((c2,d1))=\text{Enc}(0)$.
     After receiving these partial mappings, the aggregated tags are
+    • (a)
       $t_{\leq}(\{a1,b1,b2\})=\text{Enc}(1)$ and $t_{\leq}(\{c1,c2,d1\})=\text{Enc}(0)$
@@
 The tag values of $c1$, $c2$, $d1$ remain $\text{Enc}(0)$ since the initial and received tags are $\text{Enc}(0)$.
+  • 3.
     Iteration 2: The partial mappings are updated by institution B yielding
+    • (a)
       A $\rightarrow$ B: $P^{A}_{B}((a1,b1))=\text{Enc}(1),P^{A}_{B}((a1,b2))=\text{Enc}(1)$.
+    • (b)
       B $\rightarrow$ C: $P^{B}_{C}((b1,c1))=\text{Enc}(1),P^{B}_{C}((b2,c2))=\text{Enc}(1)$.
+    • (c)
       C $\rightarrow$ D: $P^{C}_{D}((c1,d1))=\text{Enc}(0),P^{C}_{D}((c2,d1))=\text{Enc}(0)$.
     After receiving these partial mappings, the aggregated tags are
-    • $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(2)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(1)$ and $t_{\leq}(d1)=\text{Enc}(0)$.
+    • (a)
+      $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(2)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(1)$, $t_{\leq}(d1)=\text{Enc}(0)$.
+  • 4.
     Iteration 3: The partial mappings are updated by B and C yielding
+    • (a)
       A $\rightarrow$ B: $P^{A}_{B}((a1,b1))=\text{Enc}(1),P^{A}_{B}((a1,b2))=\text{Enc}(1)$.
+    • (b)
       B $\rightarrow$ C: $P^{B}_{C}((b1,c1))=\text{Enc}(2),P^{B}_{C}((b2,c2))=\text{Enc}(2)$.
+    • (c)
       C $\rightarrow$ D: $P^{C}_{D}((c1,d1))=\text{Enc}(1),P^{C}_{D}((c2,d1))=\text{Enc}(1)$.
     After receiving these partial mappings, the aggregated tags are
-    • $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(3)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(3)$ and $t_{\leq}(d1)=\text{Enc}(2)$.
+    • (a)
+      $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(3)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(3)$, $t_{\leq}(d1)=\text{Enc}(2)$.
 The FinTracer tags are now sent to the FIU and decrypted. Only those accounts with a non-zero tag have received money that can be traced either directly or indirectly to suspicious NDIS payment.
+Finding key players in transaction networks
+Once the small number of accounts and entities that matched a certain ML/TF typology is detected using a technology like FinTracer, an FIU like AUSTRAC has the legal power to request much more comprehensive information about those accounts and entities from financial institutions because the request for information or notice to produce information is now targeted and backed by genuine grounds for suspicion.
+(The remaining small-scale privacy concerns can mostly be addressed using differentially private graph search algorithms like [34].) With such a dataset, it then makes sense to calculate connected components and risk metrics like PageRank scores and betweenness centrality scores, some of which have been shown to be useful features to enhance the accuracy or confidence of machine learning or rule-based financial intelligence classification models [43, 54, 62].
 Industrial grade solutions
@@
 A number of algorithms for different graph operations are included in http://github.com/AUSTRAC/ftillite/Documentation/Purgles.
-The extension of FinTracer, which allows only linear operations because of the use of additive HE, to non-linear set and graph operations to allow the detection of more-or-less arbitrarily complex criminal typologies is described in [BCF23].
-Note that differential privacy noise is added in several steps in these algorithms to avoid variations of database-reconstruction attacks [DN03, DMT07].
-Finding key players in transaction networks
-Once the small number of accounts and entities that matched a certain ML/TF typology is detected using a technology like FinTracer, an FIU like AUSTRAC has the legal power to request much more comprehensive information about those accounts and entities from financial institutions because the request for information or notice to produce information is now targeted and backed by genuine grounds for suspicion.
-(The remaining small-scale privacy concerns can mostly be addressed using differentially private graph search algorithms like [KRWY16].) With such a dataset, it then makes sense to calculate connected components and risk metrics like PageRank scores and betweenness centrality scores, some of which have been shown to be useful features to enhance the accuracy or confidence of machine learning or rule-based financial intelligence classification models [MCF+17, SvHA+19, vEDvdB+24].
+The extension of FinTracer, which allows only linear operations because of the use of additive HE, to non-linear set and graph operations to allow the detection of more-or-less arbitrarily complex criminal typologies is described in [9].
+Note that differential privacy noise is added in several steps in these algorithms to avoid variations of database-reconstruction attacks [21, 22].
+Techniques described in this section can also be extended to provide more general SQL-like queries on encrypted data, using ideas described in CryptDB [48, 63].
+In this area, Enveil and Duality provide industry-grade solutions with their products, Enveil ZeroReveal Search and Duality Query Engine.
+Such solutions allow organisations to run SQL-like queries on decentralised databases, ensuring data remains within the owner’s environment and supporting compliance with data sovereignty laws.
+Additionally, these secure systems encrypt both query content and results, keeping them private and visible only to the querier.
 4.3 Collaborative Model Learning
 Problem statement
-The use of supervised learning and anomaly detection algorithms to learn financial crime models from data has long been a topic of interest within financial intelligence [SvHA+19, vEDvdB+24], but the practical utility of such approaches has been hampered by a lack of quality labelled data in the case of supervised learning algorithms, and high false positive rates in the case of anomaly detection algorithms.
+The use of supervised learning and anomaly detection algorithms to learn financial crime models from data has long been a topic of interest within financial intelligence [54, 62], but the practical utility of such approaches has been hampered by a lack of quality labelled data in the case of supervised learning algorithms, and high false positive rates in the case of anomaly detection algorithms.
 Recent advances in privacy-preserving FL have the potential to alleviate some of these issues by allowing different business units within a financial institution and even multiple financial institutions to jointly learn models from their collective data.
@@
 All these imply that we need scalable and secure ways to perform large-scale probabilistic data matching and model learning in a privacy-preserving manner.
-The publication and reuse of machine learning models by multiple organisations also open the door to attacks like membership inference that need to be managed [MPP+21].
+The publication and reuse of machine learning models by multiple organisations also open the door to attacks like membership inference that need to be managed [44].
 Proposed solutions
@@
 Unlike transaction flow tracing, where the source and target of each transaction are uniquely identifiable for exact matching, there can be inconsistencies in how different financial institutions collect and store individual information, such as names, addresses, driver’s licences, business names, and other details.
-Consequently, data matching and entity resolution in federated learning – both within and across financial institutions – are often performed probabilistically.
-A comprehensive description of key privacy-preserving data matching/record linkage algorithms, including those that are built on Bloom Filters as described in Section 3.1, can be found in [CRS20].
+Consequently, data matching and entity resolution in federated learning - both within and across financial institutions - are often performed probabilistically.
+A comprehensive description of key privacy-preserving data matching/record linkage algorithms, including those that are built on Bloom Filters can be found in [18].
 We now describe a simplified example of federated learning to illustrate its applicability in financial intelligence.
 The task is to learn a linear regression risk model using probabilistically matched data from multiple participating parties, while preserving privacy in a distributed manner.
-The target of the regression is the individual risk score, and the features for the regression are derived from all the data held by the REs; the setup is shown in Figure 8 for three parties.
-More formally, given a training set $\{({\bf\it x}_{l},y_{l})\}_{l=1,\dots,m}$ that includes $m$ individuals, where the feature set ${\bf\it x}_{l}={\bf\it x}_{l}^{(1)}\cup\cdots\cup{\bf\it x}_{l}^{(p)}$ is the union of the features from all parties, the task is to fit a linear regression model
-${\bf\it Y}={\bf\it X}{\bf\it w}+{\bf\it\epsilon}$
-that predicts the target variable ${\bf\it Y}$ as accurate as possible.
-Figure 8: Federated learning setup for privacy-preserving risk modelling
+The target of the regression is the individual risk score, and the features for the regression are derived from all the data held by the REs.
+The setup is shown in Figure 6 for three parties.
+More formally, given a training set $\{({\bf\it x}_{l},y_{l})\}_{l=1,\dots,m}$ that includes $m$ individuals, where the feature set ${\bf\it x}_{l}={\bf\it x}_{l}^{(1)}\cup\cdots\cup{\bf\it x}_{l}^{(p)}$ is the union of the features from all parties, the task is to fit a linear regression model ${\bf\it Y}={\bf\it X}{\bf\it w}+{\bf\it\epsilon}$ that predicts the target variable $\bf\it Y$} as accurate as possible.
+Figure 6: Federated learning setup for privacy-preserving risk modelling using linear regression
 The regression training is performed using gradient descent.
@@
   1. 1.
-    The FIU generates a public-private key pair ($pk$, $sk$) and shares the public key $pk$ with the REs.
+    The FIU generates a public-private key pair ($pk$, sksk) and shares the public key $pk$ with the REs.
   2. 2.
@@
 In the remainder of this subsection, we provide a survey of some innovative use of FL that holds promise for financial intelligence.
-In the work of [SZK+22], we see a reasonably compelling demonstration of how FL can significantly enhance the detection capabilities of financial crime models.
+In the work of [59], we see a reasonably compelling demonstration of how FL can significantly enhance the detection capabilities of financial crime models.
 Through a series of case studies and rigorous experimental evaluations, this research highlights the dual achievement of improved detection performance and robust data privacy.
-The study underscores how FL can elevate the precision of financial crime detection while safeguarding sensitive information.
-Building upon this foundation, [DNM+23] introduces a novel approach that advances secure and collaborative model training across multiple financial institutions.
+Building upon this foundation, [19] introduces a novel approach that advances secure and collaborative model training across multiple financial institutions.
 By integrating HE with FL, this approach achieves secure model aggregation and parameter updates.
 The experimental results and evaluations presented reveal the approach’s effectiveness in preserving data privacy and security, thus fostering a collaborative environment for financial data analysis across institutions.
-Further extending this narrative, [HHIL+17] presents an end-to-end scheme designed for linear classifiers, employing additive HE to counteract threats from semi-honest adversaries in vertical FL scenarios.
+Further extending this narrative, [29] presents an end-to-end scheme designed for linear classifiers, employing additive HE to counter semi-honest adversaries in vertical FL scenarios.
 This research delves into the impact of entity resolution on learning performance, utilising a Taylor approximation for the loss and gradient functions.
 The application of Paillier additive HE within this context enables privacy-preserving computations, showcasing an innovative method for handling sensitive data securely.
-Additionally, [BP20] proposes a FL scheme tailored for fraud detection.
+Additionally, [13] proposes a FL scheme tailored for fraud detection.
 In this scheme, individual financial organisations enhance their model parameters with DP noise before transmitting them to a central server.
 This strategy effectively diminishes the server’s ability to recover private information, advancing the cause of privacy in fraud detection efforts.
-Though these FL techniques have yet to achieve industrial-grade maturity, they collectively offer a glimpse into the forefront of financial intelligence.
+The above FL techniques collectively offer a glimpse into the forefront of financial intelligence.
 Their contributions highlight ongoing advancements and set the stage for future innovations in the field.
 Industrial grade solutions
-There are several industrial-grade solutions like [ZCN17, ZNCC18] for conducting probabilistic matching of databases with millions to billions of entities and these can be used in conjunction with techniques like Bloom filters and HE to produce somewhat scalable privacy-preserving probabilistic data-matching algorithms.
-While federated learning is now a mature technology, progress in using federated learning at the industrial scale for financial intelligence applications has been slow because, in most cases, there is not an abundance of labelled data to support the acquisition of robust and trusted machine-learning models with low false-positive rates.
-With further advances in FL and a general uplift in the maturity of FIUs in adopting privacy technologies, we expect to see more progress in this area in the future.
-5 Discussion and Conclusion
+There are several industrial-grade solutions like [64, 65] for conducting probabilistic matching of databases with millions to billions of entities and these can be used in conjunction with techniques like Bloom filters and HE to produce somewhat scalable privacy-preserving probabilistic data-matching algorithms.
+Privitar SecureLink [^7] SecureLink website offers a solution for securely joining data from various contributors without revealing sensitive identifiers or directly linking data across organisations.
+While federated learning is now a mature technology, progress in using federated learning at the industrial scale for financial intelligence applications is still somewhat hampered by general lack of labelled data to support the acquisition of robust and trusted machine-learning models with low false-positive rates.
+Having said that, several case studies and investigations have validated and implemented privacy-preserving solutions for specific scenarios.
+In particular, industrial solutions developed by Enveil, Duality, and Inpher have scaled effectively to address real-world challenges, with some initiatives leading to collaborations with other organisations to create domain specific products.
+The details are described in Section 5.
+5 Real-World Case Studies
+In this section, we describe several case studies that have been trialled and, in some cases, deployed in real-world applications.
+The presentation provides an update on some key projects first listed in [41] and includes more recent additional work; the emphasis is on the business problem, participating financial institutions, and achieved business outcomes (when that is available.) The case studies vary in scope and complexity, with some focusing on a singular use case while others cover all the three major use cases described in Section 4.
+It is important to note that in the practical context, AML/CTF use cases can be affected by a large number of factors, including computational cost, operational cost, technical complexity, efficiency, data availability, quality and interoperability.
+These case studies are summarised in Table 1.
+Table 1: Overview of several case studies by different jurisdictions.
+Project Participants Use case
+Enveil ZeroReveal Inter-Bank Secure and Private Data Collaboration [41] An EU-based financial institution and a third-party data provider Transaction flow tracing and typology matching
+Duality SecurePlus Query [41] UK and Canadian Banks Transaction flow tracing and typology matching
+DANIE consortium [41] UK-based consortium of banks Suspicious persons list
+Privitar SecureLink [41] Public and private Institutes in an EU member state Suspicious persons list
+BNY Mellon and Inpher XOR [31] BNY Mellon Collaborative learning
+UK Tri-bank initiative [42] Three UK banks Transaction flow tracing
+AUSTRAC’s Fintel Alliance Alerting Project [14, 58] Led by AUSTRAC Australia and delivered by the public-private partnership Fintel Alliance Transaction flow tracing
+AML Bridge Platform [25] Ten banks and several non-banks in Estonia Transaction flow tracing and typology matching
+Transaction Monitoring Netherlands (TMNL) [61] Five banks in Netherlands Transaction flow tracing
+DeepProtect [32] Five banks in Japan Collaborative learning
+Project Aurora [8] Led by Bank for International Settlements and participated by 29 financial institutions in six European countries Collaborative learning
+COSMIC [25, 1] Owned and operated by the Monetary Authority of Singapore (MAS).
+The is a co-development with the Commercial Affairs Department (CAD) of the Singapore Police Force, and six major banks (Citibank, DBS, HSBC, OCBC, Standard Chartered Bank and UOB).
+Transaction flow tracing and typology matching
+Enveil ZeroReveal Inter-Bank Private Collaboration [41]
+This is an initiative, started in 2019, by the UK Financial Conduct Authority (FCA) Global AML and Financial Crime TechSprint and Enveil to develop a system that could enable data and knowledge sharing between stakeholders to identify and prevent money flows within and between criminal networks.
+This initiative was supported by several private UK banks including EY, BAE Systems, Refinitiv, HSBC, Barclays and ING.
+Their proposed system used homomorphic encryption to provide secure and automated multiparty collaboration which allows banks to query about higher risk customers.
+Since that FCA TechSprint, the Enveil suite of solutions have developed in maturity and are now used in multiple sensitive domains.
+[^8] Enveil Wins Army Linchpin Contract to Deliver Secure AI
+Duality SecurePlusTM Query on Financial Data [41]
+The product Duality SecurePlusTM Query enables financial crime and compliance teams to carry out collaborative investigations across groups, institutions and borders by keeping sensitive data encrypted throughout the process.
+This product allows cross-border collaboration between financial crime and compliance teams.
+The product allows participant to query distributed data silos while keeping the sensitive data encrypted throughout the querying process.
+A case study between several banks in the UK and Canada demonstrated the capability of banks to match customers that includes external data sets, without disclosing their original query to the data owner.
+This allows different financial intelligence teams to securely query each other’s data.
+The system also supports to reduce false positives, improve outcomes, increase efficiency, and most importantly effectively expose financial crime networks.
+[^9] Leverage Secure Data Sharing to Fight Money Laundering
+DANIE Consortium [41]
+Project DANIE is a collaborative initiative involving multiple financial institutions from EU Member States, aimed at enhancing banking data quality through the utilisation of privacy technologies.
+The project employs a combination of secure hardware and software components, along with secure multiparty computation and homomorphic encryption.
+The primary objective of DANIE is to improve data quality by matching datasets across participating organisations to identify discrepancies and outliers, while simultaneously preserving the anonymity and security of the matched data through these technologies.
+In addition to its application in enhancing data reconciliation and subsequently improving the Know Your Customer (KYC) process, DANIE has the potential to be adapted for addressing the Suspicious Person List (SPL) cases discussed in Section 4.1.
+This can be achieved by encrypting an SPL and uploading it to the DANIE platform to request confirmation of each individual’s presence within the participating financial institutions.
+BNY Mellon and Inpher XOR [31]
+The Bank of New York Mellon Corporation uses machine-learning (ML) based fraud detection algorithms to find hidden patterns in money laundering networks.
+Together with the Inpher’s XOR platform, BNY Mellon built a novel collaborative fraud detection framework leveraging Secure Multiparty Computation.
+Inpher’s technology enabled BNY Mellon to access sensitive data across borders or jurisdictional silos for training ML algorithms such as linear and logistic regression as well as XGBoost.
+The solution achieved a high level of precision, computing models identically as if they were computed in clear text.
+BNY Mellon’s project demonstrated that the number of false positives was reduced by 20% and the number of true positives was increased by 20%.
+The solution thus reduced investigation workload and increased fraud detection efficiency.
+UK Tri-bank initiative [42]
+This is a proof-of-concept (POC) system developed by three UK banks to analyse historical data on small and medium sized corporate client datasets.
+The system used the FutureFlow Cross-Bank Financial Crime Analytics Platform to analyse transaction data between banks.
+In their POC study, banks analysed approximately 200,000 accounts for money laundering risk.
+In this system, the transaction data is encrypted to protect sensitive data before using to identify links between the transactions.
+This allowed the banks to identify potential money laundering networks based on the patterns in the transaction data.
+AUSTRAC’s Fintel Alliance Alerting Project [14]
+The FINTEL Alliance comprises representatives from both public and private partner organizations that collaborate to address complex or emerging crimes affecting the community, necessitating a joint public-private approach.
+The aim of this project is to develop different approaches which would allow the identification of complex money transferring networks used by criminal gangs.
+The project used a federated platform with homomorphic encryption technologies where entities could report or provide access to the agreed dataset via an API.
+The project was tested in a simulated federated architecture using a sample of real data, covering more than 100 million accounts.
+AML Bridge Platform [25]
+AML Bridge is a secure platform provided by Salv.
+The platform uses an end-to end encryption technique which allows financial institutions to securely exchange of pseudonymised transaction data.
+The platform was first tested as a POC on four of Estonia’s largest banks and later expanded to all Estonian banks.
+In their POC data owners can also limit what queries it allows from its own data sources while the requesting party is not revealed to the data owner.
+In their querying protocol, the platform first encrypts the data at the source.
+Next data is aggregated with other parties’ and stored in three trusted hosts using secure multiparty computation.
+Finally, within the requesting party’s trusted environment, the platform aggregates the data results without revealing the original data source(s).
+Transaction Monitoring Netherlands (TMNL) [61]
+This is an initiative consists of five Dutch banks including ABN AMRO, ING, Rabobank, Triodos Bank and De Volksbank.
+The main aim of this initiative is to combine transaction data of business accounts from these banks to gain a complete view of money flows by criminal networks.
+The initiative uses a central trusted data store where each bank could upload their data.
+The data could then aggregate by combining these transactions and the consortium (the group of banks) could monitor and report potentially suspicious patterns to the others.
+central data store where the consortium combines these transactions, monitors them, and reports potentially suspicious patterns.
+The central data store encrypts data at rest, which can only be viewed when when suspicions of money laundering or terrorist financing arise Sensitive information is encrypted and viewed only when suspicions of money laundering or terrorist financing arise.
+DeepProtect [32]
+This is research project by National Institute of Information and Communications Technology (NICT) in Japan.
+In this project, NICT collaborated with five banks in Japan including Chiba Bank, Ltd., MUFG Bank, Ltd., the Chugoku Bank, Ltd., Sumitomo Mitsui Trust Bank, Ltd., and the Iyo Bank, Ltd.
+The aim of this project was to build models for detecting criminals’ bank accounts and fraudulent transactions in customers accounts.
+DeepProtect employs a FL framework which enables each bank to run a local model using its own data while collaboratively train a global model with data from other banks.
+For privacy preservation, the implementation used a homomorphic encryption scheme in a symmetric key environment.
+Project Aurora [8]
+This is an initiative led by the Bank for International Settlements (BIS).
+The aim of this project is to identify money laundering between banks and cross-border financial institutions by analysing the connected payment data.
+The project uses different privacy and security techniques with ML and network analysis methods for identifying potential suspicious transactions.
+Phase 1 of the project demonstrated that the combinations of these technologies provides more effective identification of money laundering networks compared to traditional siloed and rules-based approaches.
+In their experimental setting, consisting 29 financial institutions in six different European countries, the project detected potentially up to 3 times more money laundering involved complex schemes and reduced false positives by up to 80%.
+COSMIC [1]
+COSMIC is a secure digital platform co-developed by Monetary Authority of Singapore (MAS) and six major commercial banks in Singapore including DBS, OCBC, UOB, Citibank, HSBC and Standard Chartered.
+Released in 2024, COSMIC allows financial institutions to securely share transaction data and information about suspicious accounts with one another.
+The initial aim of COSMIC is to provide a easier and efficient data sharing platform for FIs to detect and thereby deter criminal activity.
+Currently, information sharing in COSMIC is voluntary.
+The platform currently supports three key financial crime risks, (1) misuse of legal persons, (2) misuse of trade finance for illicit purposes and (3) proliferation financing.
+COSMIC is owned and operated by the MAS.
+The above shortlist of projects is quite compelling and is demonstrative of the increasing global interest in the adoption of maturing privacy technologies by major financial institutions, both public and private, to combat financial crimes.
+6 Discussion
+Legislative Reform and Privacy Technologies
 Financial intelligence equips law enforcement and financial institutions with sophisticated analytical tools and data-driven methodologies to proactively detect and prevent financial crimes.
@@
 While public and private sectors, industry associations and regulatory bodies are actively working to harmonise AML/CTF and data privacy regulations, such changes often involve lengthy and uncertain political approval and implementation processes.
-From the practical perspective of being able to do something about complex financial crimes today, we advocate for the integration of privacy-preserving technologies within financial intelligence tools, not just as a mitigation strategy while waiting for legal experts to navigate the complex AML/CTF reform landscape [Bug18], but to inform those AML/CTF reform discussions.
+From the practical perspective of being able to do something about complex financial crimes today, we advocate for the integration of privacy-preserving technologies within financial intelligence tools, not just as a mitigation strategy while waiting for legal experts to navigate the complex AML/CTF reform landscape [11], but also to inform those AML/CTF reform discussions.
 This approach offers several benefits.
@@
 Second, adopting privacy-preserving methods helps institutions build trust with customers and stakeholders by demonstrating a commitment to both data protection and effective financial crime prevention.
-Third, privacy-preserving solutions offer a flexible framework that can adapt to evolving regulatory requirements, helping institutions stay compliant as laws and regulations are updated.
+Finally, privacy-preserving solutions offer a flexible framework that can adapt to evolving regulatory requirements, helping institutions stay compliant as laws and regulations are updated to address new financial technologies and emerging criminal threats.
+Trade-offs and Practical Considerations
+In practical financial intelligence applications, selecting the most appropriate privacy techniques is crucial for ensuring data security, regulatory compliance, and maintaining user trust at sustainable costs.
+There are often trade-offs involved in such choices.
+With the increasing industry focus on responsible AI practices [20, 33], there is also an emerging need to understand the possible interactions between privacy requirements and related issues like explainability and fairness [5, 49, 60].
+Scholars and practitioners must look ahead to how AI itself may transform AML/CTF, potentially enabling new, automated, or hybrid methods especially when combined with emerging technologies such as decentralised autonomous organisations [30].
+Strategic foresight and scenario-based methodologies may therefore be essential to anticipate and prepare for the future evolution of AI-enabled AML/CTF.
+To inform those broader issues, we offer here a few observations on how privacy-utility trade-offs are usually handled for the different privacy technologies.
+First of all, it is worth noting that the use of homomorphic encryption (HE) and secure multiparty computation (MPC) in financial-intelligence algorithms like private set intersection, graph pattern matching, and statistical model learning do not usually result in any change to the output of those algorithms.
+This is because HE and MPC allow an algorithm to be evaluated in an encrypted space but they do not otherwise change the nature of a given algorithm.
+The primary consideration when using HE and MPC is thus whether the specific algorithm and parameters chosen provide, at reasonable organisational and monetary costs, sufficient provable security against adversaries with access to potentially large computational resources and arbitrary background and side information.
+In contrast, the use of differential privacy (DP) and federated learning (FL) to privatise a financial-intelligence algorithm can result in actual changes to the output of the algorithm.
+There are rich and well-established frameworks [23, 35, 47] to study trade-offs between privacy and utility when using DP, and these frameworks work well in conjunction with standard techniques for model evaluation and selection [51] in both federated learning and more typical setups.
+Addressing these questions will be central to future research in this field.
+7 Conclusion
 We hope this paper enhances public understanding of privacy-preserving financial intelligence through the various AML/CTF use cases we explore.
@@
 Having said that, for financial institutions and regulatory bodies, adopting privacy technologies involves careful consideration of both technological and operational aspects.
-Effective implementation strategies must balance the trade offs between data privacy and the quality and efficiency of analysis.
+Effective implementation strategies must balance the trade-offs between data privacy and the quality and efficiency of analysis.
 While these technologies address immediate challenges, they are not a substitute for comprehensive regulatory reforms.
-Continuous refinement of regulations is essential to align financial crime prevention efforts with data privacy goals.
+Further and continuous refinement of regulations is essential to align financial crime prevention efforts with data privacy goals.
 Acknowledgements
-We would like to thank the many former colleagues in AUSTRAC and financial crime units in major Australian financial institutions who spent many hours teaching us the intricacies of the financial intelligence business.
-We are also grateful to many of them for continuing to play an important role to keep our financial system and our country safe.
+We thank the many former colleagues in AUSTRAC and financial crime units in major Australian financial institutions who spent many hours teaching us the intricacies of the financial intelligence business.
+We are grateful to many of them for continuing to play an important role to keep our financial system and our country safe.